### v1

Abstract
In this paper, we introduce the applications of third-order reduced biquaternion tensors in color video processing. We first develop algorithms for computing the singular value decomposition (SVD) of a third-order reduced biquaternion tensor via a new Ht-product. As theoretical applications, we define the Moore-Penrose inverse of a third-order reduced biquaternion tensor and develop its characterizations. In addition, we discuss the general (or Hermitian) solutions to reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$ as well as its least-square solution. Finally, we compress the color video by this SVD, and the experimental data shows that our method is faster than the compared scheme.
Key words: Reduced biquaternion tensor; Ht-product; Singular value decomposition; Moore-Penrose inverse.
AMS 2010 Subject Classification: 15A09, 15A18, 15A23, 15A24, 15B33, 39B42.
1 Introduction
In 1843, Hamilton extended the complex number field $\mathbb{C}$ to quaternions which is a four dimensional division algebra over the real number field $\mathbb{R}$. It is well-known that Hamilton quaternions have successful applications in signal and color image processing (e.g., [5]). To keep the commutativity of usual number systems, in 1892, Segre ([18]) discovered the following extension:
$\mathbb{H}_{c}=\{a_{1}+a_{2}i+a_{3}j+a_{4}k\ |\ i^{2}=-1,\ j^{2}=1,\ ij=ji=k,\ a_{1},\dots,a_{4}\in\mathbb{R}\}.$
Clearly, $k^{2}=-1,\ ki=ik=-j,\ jk=kj=i$. $\mathbb{H}_{c}$ is a four dimensional commutative algebra over $\mathbb{R}$ and contains zero divisors. It is called Segre biquaternion ([14]), commutative quaternion ([9, 12, 13]) or reduced biquaternion ([13, 17]). In this paper, we call it “reduced biquaternion” ( or RB for short).
Reduced biquaternion $\mathbb{H}_{c}$ has numerous applications in many areas. In signal and color image processing, Pei et al. ([12, 13]) show that the operations of the discrete reduced biquaternion Fourier transforms and their corresponding convolutions and correlations are much simpler than the existing implementations of the Hamilton quaternions, and many questions can be performed simultaneously by $\mathbb{H}_{c}$. They use reduced biquaternion matrices to represent the color images, then compress the images by the SVDs of reduced biquaternion matrices (RBSVDs). For comparison, they found that the calculating of RBSVD is faster than SVD of a quaternion matrix (QSVD), and the reconstruction quality is also better than one by QSVD method.
By adopting the special structure of a real representation of a reduced biquaternion matrix, Ding et al. [4] investigated the least-squares special minimal norm solutions to RB matrix equation $AX=B$ and applied the least-squares minimal norm RB solution to the color image restoration.
A tensor can be regarded as a multidimensional array. People use tensors to collect and employ multi-way data. Tensors and tensor decompositions ([3, 6, 7, 8, 10, 15, 19]) have been applied in many areas such as signal processing, computer science, data mining, neuroscience, and more. Studying tensor rank decomposition may go back to the 1960s: the CANDECOMP/PARAFAC (CP) decomposition ([3]) and the Tucker decomposition ([19]). In [6], Kilmer and Martin defined a closed multiplication operation (called t-product) for two real tensors, and considered some important theoretical properties and practical tensor approximation problems. Qin et al. [15] defined a tensor product for third-order quaternion tensors and discussed the Qt-SVD of a third-order quaternion tensor with applications in color video.
Motivated by the works mentioned above, in particular, the applications of reduced biquaternion matrices in color image processing, we introduce the applications of reduced biquaternion tensors in color video compression.
This paper is organized as follows. We first introduce some operations over reduced biquaternion tensors, and then we define the Ht-product and explore some important properties of third-order RB tensors in Section 2. In Section 3, we investigate the Ht-SVD of a third-order RB tensor and develop a method to compute it. Two theoretical applications of Ht-SVD are given in Sections 4 and 5. That is, in Section 4, the Moore-Penrose inverse of a third-order RB tensor is defined and some properties including algebraic expression derived by the Ht-SVD are discussed. In Section 5, we consider the general/Hertmian/least-squares solutions to reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$. Finally, in Section 6, we apply this Ht-SVD in color video compression. The experimental data shows that our method has an excellent performance comparing with another method.
2 The Ht-product of third-order tensors over $\mathbb{H}_{c}$
Throughout this paper, Euler script letters $(\mathcal{A,B,\cdots})$ are used to refer to tensors, while the capital letters $(A,B,\cdots)$ represent matrices. The notations $R^{n_{1}\times n_{2}}$ and $R^{n_{1}\times n_{2}\times n_{3}}$ represent the set of all the matrices of dimension $n_{1}\times n_{2}$ and all the third-order tensors of dimension $n_{1}\times n_{2}\times n_{3}$ over algebra $R$ respectively. For a third-order tensor $\mathcal{A}$, we denote the $i$th horizontal, the $j$th lateral and the $k$th frontal slice by $\mathcal{A}(i,:,:),\mathcal{A}(:,j,:)$ and $\mathcal{A}(:,:,k)$ respectively. For simplicity, let $A^{(k)}$ represent the $k$th frontal slice $\mathcal{A}(:,:,k)$.
Recall that an element $q\in\mathbb{H}_{c}$ can be expressed as $q=q_{a}+jq_{b}=q_{(c),1}e_{1}+q_{(c),2}e_{2}$, where $q_{a}=q_{0}+q_{1}i,q_{b}=q_{2}+q_{3}i,$ $q_{(c),1}=q_{a}+q_{b},\ q_{(c),2}=q_{a}-q_{b}\in\mathbb{C}$, and $e_{1}=(1+j)/2,e_{2}=(1-j)/2\in\mathbb{H}_{c}$. Clearly, $e_{1}$ and $e_{2}$ are two special reduced biquaternions satisfying $e_{1}=e_{1}^{2}=\cdots=e_{1}^{n},e_{2}=e_{2}^{2}=\cdots=e_{2}^{n}$ and $e_{1}e_{2}=0$ ([13]).
For any matrix $A\in\mathbb{H}_{c}^{m\times n}$, we can write it in two forms, that is, $A=A_{0}+A_{1}i+A_{2}j+A_{3}k\doteq B_{1}+jB_{2}$ with $A_{0},\dots,A_{3}\in\mathbb{R}^{m\times n},B_{1}=A_{0}+A_{1}i,B_{2}=A_{2}+A_{3}i\in\mathbb{C}^{m\times n}$. Then its conjugate transpose is given by $A^{\ast}=A_{0}^{T}-A_{1}^{T}i+A_{2}^{T}j-A_{3}^{T}k=B_{1}^{\ast}+jB_{2}^{\ast}$. Moreover, $A\in\mathbb{H}^{n\times n}_{c}$ is said to be Hermitian (resp. unitary) if and only if $A=A^{\ast}$ (resp. $AA^{\ast}=A^{\ast}A=I_{n}$).
Furthermore, if we rewrite a $A\in\mathbb{H}_{c}^{m\times n}$ as $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}$, where $A_{(c),1}$ and $A_{(c),2}$ are complex matrices, then the conjugate transpose of $A$ is $A^{\ast}=A_{(c),1}^{\ast}e_{1}+A_{(c),2}^{\ast}e_{2}$. Thus for any $B=B_{(c),1}e_{1}+B_{(c),2}e_{2}\in\mathbb{H}_{c}^{n\times l}$, $AB=A_{(c),1}B_{(c),1}e_{1}+A_{(c),2}B_{(c),2}e_{2}$, and we can use it to prove that $(AB)^{\ast}=B^{\ast}A^{\ast}$.
The block circulant matrix $\textbf{circ}(\mathcal{A})$ generated by a third-order tensor $\mathcal{A}$’s frontal slices $A^{(1)},$ $A^{(2)},\cdots,A^{(n_{3})}$ is given as
$\textbf{circ}(\mathcal{A})=\begin{bmatrix}A^{(1)}&A^{(n_{3})}&\cdots&A^{(2)}\\ A^{(2)}&A^{(1)}&\cdots&A^{(3)}\\ \vdots&\vdots&\ddots&\vdots\\ A^{(n_{3})}&A^{(n_{3}-1)}&\cdots&A^{(1)}\\ \end{bmatrix}.$
We will use Vec operator to convert a tensors $\mathcal{A}$ into a block matrix $\textbf{Vec}(\mathcal{A})$:
$\textbf{Vec}(\mathcal{A})=\begin{bmatrix}A^{(1)}\\ A^{(2)}\\ \vdots\\ A^{(n_{3})}\\ \end{bmatrix},$
and the inverse operation Fold to take the block matrix $\textbf{Vec}(\mathcal{A})$ back to a tensor:
$\textbf{Fold}(\textbf{Vec}(\mathcal{A}))=\mathcal{A}.$
Moreover, we use the frontal slices of $\mathcal{A}$ to define a block diagonal matrix as following:
$\textbf{diag}({\mathcal{A}})\doteq\begin{bmatrix}{A}^{(1)}&&&\\ &{A}^{(2)}&&\\ &&\ddots&\\ &&&{A}^{(n_{3})}\\ \end{bmatrix}.$ (1)
Note that by using the normalized Discrete Fourier transformation (DFT) matrix, a complex circulant matrix can be diagonalized ([6], [15]). Suppose $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ is a third-order complex tensor. Then
$(F_{n_{3}}\otimes I_{n_{1}})\textbf{circ}(\mathcal{A})(F_{n_{3}}^{\ast}\otimes I_{n_{2}})=\begin{bmatrix}\widehat{A}^{(1)}&&&\\ &\widehat{A}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{A}^{(n_{3})}\\ \end{bmatrix}=\textbf{diag}(\widehat{\mathcal{A}})\in\mathbb{C}^{n_{3}n_{1}\times n_{3}n_{2}},$ (2)
where $\otimes$ denotes the Kronecker product, $\widehat{A}^{(i)}$ ($i=1,\cdots,n_{3}$) are the frontal slices of the tensor $\widehat{\mathcal{A}}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ which is the result of DFT of $\mathcal{A}$ along the third mode, $F_{n_{3}}$ is unitary and $F_{n_{3}}(i,j)=\frac{1}{\sqrt{n_{3}}}\omega^{(i-1)(j-1)},\ \omega=exp(-\frac{2\pi i}{n_{3}})$.
Definition 2.1
The DFT of $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ along the third mode is a tensor $\mathcal{\widehat{A}}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ satisfying
$\widehat{\mathcal{A}}(i,j,:)=\sqrt{n_{3}}F_{n_{3}}\mathcal{A}(i,j,:).$
We can verify the relation between $\mathcal{A}$ and $\widehat{\mathcal{A}}$ as follows.
$\textbf{Vec}(\widehat{\mathcal{A}})=\begin{bmatrix}\widehat{A}^{(1)}\\ \widehat{A}^{(2)}\\ \vdots\\ \widehat{A}^{(n_{3})}\\ \end{bmatrix}=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}).$ (3)
Now we extend $t$-product of two third-order real tensors defined in [6] to reduced biquaternions.
Definition 2.2
(Ht-product) Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$. Then $Ht$-product of $\mathcal{A}$ and $\mathcal{B}$ is defined as
$\mathcal{A}\ast_{Ht}\mathcal{B}=\textbf{Fold}((\textbf{circ}(\mathcal{A}_{1,i})+j\textbf{circ}(\mathcal{A}_{j,k}))\textbf{Vec}(\mathcal{B}))\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}},$
where $\mathcal{A}=\mathcal{A}_{1,i}+j\mathcal{A}_{j,k}$, $\mathcal{A}_{1,i},\mathcal{A}_{j,k}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}.$
The addition and scalar multiplication rules of tensors over $\mathbb{H}_{c}$ are defined in a usual way.
Theorem 2.3
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}},\ \mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$ and $\mathcal{C}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}$ with the corresponding DFTs $-^\widehat{\mathcal{A}},\ \widehat{\mathcal{B}}$ and $\widehat{\mathcal{C}}$. Then $\mathcal{A}\ast_{Ht}\mathcal{B}=\mathcal{C}$ if and only if $\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}})=\textbf{diag}(\widehat{\mathcal{C}}).$
Proof.  Since $\widehat{\mathcal{A}}$ is the DFT of $\mathcal{A}=\mathcal{A}_{1,i}+j\mathcal{A}_{j,k}$ along the third mode, it follows from (3) and $F_{n_{3}}j=jF_{n_{3}}$ that
$\textbf{Vec}(\widehat{\mathcal{A}})$ $=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})(\textbf{Vec}(\mathcal{A}_{1,i})+j\textbf{Vec}(\mathcal{A}_{j,k}))$
$=\textbf{Vec}(\widehat{\mathcal{A}}_{1,i})+j\textbf{Vec}(\widehat{\mathcal{A}}_{j,k}),$
which implies
$\textbf{diag}(\widehat{\mathcal{A}})=\textbf{diag}(\widehat{\mathcal{A}}_{1,i})+j\textbf{diag}(\widehat{\mathcal{A}}_{j,k}).$ (4)
By Definition 2.2 and identities (2) and (3), we have
$\textbf{Vec}(\mathcal{C})$ $=\textbf{Vec}(\mathcal{A}\ast_{Ht}\mathcal{B})$
$=(\textbf{circ}(\mathcal{A}_{1,i})+j\textbf{circ}(\mathcal{A}_{j,k}))\textbf{Vec}(\mathcal{B})$
$=(F_{n_{3}}^{\ast}\otimes I_{n_{1}})[\textbf{diag}(\widehat{\mathcal{A}}_{1,i})+j\textbf{diag}(\widehat{\mathcal{A}}_{j,k})](F_{n_{3}}\otimes I_{n_{2}})\textbf{Vec}(\mathcal{B})$
$=\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})[\textbf{diag}(\widehat{\mathcal{A}}_{1,i})+j\textbf{diag}(\widehat{\mathcal{A}}_{j,k})]\textbf{Vec}(\widehat{\mathcal{B}}).$
Thus
$\textbf{Vec}(\widehat{\mathcal{C}})=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{C})=[\textbf{diag}(\widehat{\mathcal{A}}_{1,i})+j\textbf{diag}(\widehat{\mathcal{A}}_{j,k})]\textbf{Vec}(\widehat{\mathcal{B}}).$
Combing with equation (4) yields $\textbf{Vec}(\widehat{\mathcal{C}})=\textbf{diag}(\widehat{\mathcal{A}})\textbf{Vec}(\widehat{\mathcal{B}})$. Therefore,
$\textbf{diag}(\widehat{\mathcal{C}})=\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}}).$
$\blacksquare$
According to the above result, we can give an equivalent definition of the Ht-product of two third-order reduced biquaternion tensors.
Definition 2.4
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$. The $Ht$-product of $\mathcal{A}$ and $\mathcal{B}$ is defined as
$\mathcal{A}\ast_{Ht}\mathcal{B}={\textbf{Fold}}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}})(e\otimes I_{n_{4}})\right),$
where $e$ is an $n_{3}\times 1$ vector with all the entries equal to 1.
This Ht-product obeys some usual algebraic operation laws, such as the associative law and the distributive law.
Proposition 2.5
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}},\ \mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$ and $\mathcal{C}\in\mathbb{H}_{c}^{n_{4}\times n_{5}\times n_{3}}$. Then
$(\mathcal{A}\ast_{Ht}\mathcal{B})\ast_{Ht}\mathcal{C}=\mathcal{A}\ast_{Ht}(\mathcal{B}\ast_{Ht}\mathcal{C}).$
Proof.  By Definition 2.4, we have
$(\mathcal{A}\ast_{Ht}\mathcal{B})\ast_{Ht}\mathcal{C}=$ $\textbf{Fold}(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}})(e\otimes I_{n_{4}}))\ast_{Ht}\mathcal{C}$
$=$ $\textbf{Fold}(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})(\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}}))\textbf{diag}(\widehat{\mathcal{C}})(e\otimes I_{n_{5}})),$
and
$\mathcal{A}\ast_{Ht}(\mathcal{B}\ast_{Ht}\mathcal{C})=$ $\mathcal{A}\ast_{Ht}\textbf{Fold}(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{2}})\textbf{diag}(\widehat{\mathcal{B}})\textbf{diag}(\widehat{\mathcal{C}})(e\otimes I_{n_{5}}))$
$=$ $\textbf{Fold}(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})\textbf{diag}(\widehat{\mathcal{A}})(\textbf{diag}(\widehat{\mathcal{B}})\textbf{diag}(\widehat{\mathcal{C}}))(e\otimes I_{n_{5}})).$
Notice that $\textbf{diag}(\widehat{\mathcal{A}}),\textbf{diag}(\widehat{\mathcal{B}})$ and $\textbf{diag}(\widehat{\mathcal{C}})$ are matrices over $\mathbb{H}_{c}$. Hence the associative law holds, and thus the result holds. $\blacksquare$
In a similar way as in Proposition 2.5, we have the following distributive laws. For brevity, we omit the proofs.
Proposition 2.6
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}},\ \mathcal{B},\ \mathcal{C}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$ and $\mathcal{D}\in\mathbb{H}_{c}^{n_{4}\times n_{5}\times n_{3}}$. Then
$\mathcal{A}\ast_{Ht}(\mathcal{B}+\mathcal{C})=\mathcal{A}\ast_{Ht}\mathcal{B}+\mathcal{A}\ast_{Ht}\mathcal{C},$
$(\mathcal{B}+\mathcal{C})\ast_{Ht}\mathcal{D}=\mathcal{B}\ast_{Ht}\mathcal{D}+\mathcal{C}\ast_{Ht}\mathcal{D}.$
Suppose $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$. Using the method in [15], we calculate the conjugate transpose $\mathcal{A}^{\ast}\in\mathbb{C}^{n_{2}\times n_{1}\times n_{3}}$ of $\mathcal{A}$ as follows: first conjugately transposing each frontal slice of $\mathcal{A}$ first, and reversing the order of conjugately transposed frontal slices $2$ through $n_{3}$ next.
Definition 2.7
The conjugate transpose $\mathcal{A}^{\ast}$ of $\mathcal{A}=\mathcal{A}_{1,i}+j\mathcal{A}_{j,k}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ with $\mathcal{A}_{1,i},\mathcal{A}_{j,k}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ is defined as $\mathcal{A}^{\ast}=\mathcal{A}_{1,i}^{\ast}+j\mathcal{A}_{j,k}^{\ast}$ with $\mathcal{A}_{1,i}^{\ast},\mathcal{A}_{j,k}^{\ast}\in\mathbb{C}^{n_{2}\times n_{1}\times n_{3}}$.
It is not difficult to verify that $(\mathcal{A}^{\ast})^{\ast}=\mathcal{A}$ and $(\mathcal{A}+\mathcal{B})^{\ast}=\mathcal{A}^{\ast}+\mathcal{B}^{\ast}$ for $\mathcal{A},\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$.
Definition 2.8
The identity tensor $\mathcal{I}_{nnm}\in\mathbb{H}_{c}^{n\times n\times m}$ (simply denoted by $\mathcal{I}$) is the tensor in which the first frontal slice is $I_{n}$ and the other slice matrices are zero.
We can obtain that $\mathcal{I}\ast_{Ht}\mathcal{I}=\mathcal{I}$$\mathcal{I}^{\ast}=\mathcal{I}$ and $\mathcal{A}\ast_{Ht}\mathcal{I}=\mathcal{I}\ast_{Ht}\mathcal{A}=\mathcal{A}.$
Definition 2.9
A tensor $\mathcal{A}\in\mathbb{H}_{c}^{n\times n\times m}$ is called invertible if there is a tensor $\mathcal{B}\in\mathbb{H}_{c}^{n\times n\times m}$ such that $\mathcal{A}\ast_{Ht}\mathcal{B}=\mathcal{I}$ and $\mathcal{B}\ast_{Ht}\mathcal{A}=\mathcal{I}$.
By using Proposition 2.5 and the fact $\mathcal{A}\ast_{Ht}\mathcal{I}=\mathcal{I}\ast_{Ht}\mathcal{A}=\mathcal{A}$, we have the following result.
Proposition 2.10
The inverse of a Segre quaternion tensor $\mathcal{A}\in\mathbb{H}_{c}^{n\times n\times m}$ is unique if it exists, denoted by $\mathcal{B}=\mathcal{A}^{-1}$.
Definition 2.11
Let $\mathcal{A}\in\mathbb{H}_{c}^{n\times n\times m}$. $\mathcal{A}$ is called idempotent if $\mathcal{A}\ast_{Ht}\mathcal{A}=\mathcal{A}$; $\mathcal{A}$ is called unitary (resp. Hermitian) if $\mathcal{A}\ast_{Ht}{\mathcal{A}}^{\ast}={\mathcal{A}}^{\ast}\ast_{Ht}\mathcal{A}=\mathcal{I}\ (resp.\ \mathcal{A}={\mathcal{A}}^{\ast}).$
Lemma 2.12
Let $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$. Then $\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$
Proof.  It is easy to see that $\textbf{circ}({\mathcal{A}}^{\ast})=(\textbf{circ}({\mathcal{A}}))^{\ast}$ for $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}.$ According to the equality (2), we have
$\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})$ $=(F_{n_{3}}\otimes I_{n_{2}})\textbf{circ}({\mathcal{A}}^{\ast})(F_{n_{3}}^{\ast}\otimes I_{n_{1}})$
$=(F_{n_{3}}\otimes I_{n_{2}})(\textbf{circ}({\mathcal{A}}))^{\ast}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})$
$=(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$
$\blacksquare$
The above result over $\mathbb{C}$ can be extended to $\mathbb{H}_{c}$ as follows.
Proposition 2.13
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then $\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$
Proof.  Since $\widehat{{\mathcal{A}}^{\ast}}$ is the DFT of ${\mathcal{A}}^{\ast}$ and $F_{n_{3}}j=jF_{n_{3}}$, it follows from (3) that
$\textbf{Vec}(\widehat{{\mathcal{A}}^{\ast}})$ $=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{2}})\textbf{Vec}({\mathcal{A}}^{\ast})$
$=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{2}})(\textbf{Vec}({\mathcal{A}}_{1,i}^{\ast})+j\textbf{Vec}({\mathcal{A}}_{j,k}^{\ast}))$
$=\textbf{Vec}(\widehat{{\mathcal{A}}_{1,i}^{\ast}})+j\textbf{Vec}(\widehat{{\mathcal{A}}_{j,k}^{\ast}}),$
which is equivalent to
$\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=\textbf{diag}(\widehat{{\mathcal{A}}_{1,i}^{\ast}})+j\textbf{diag}(\widehat{{\mathcal{A}}_{j,k}^{\ast}}).$ (5)
Similarly, we have $\textbf{diag}(\widehat{\mathcal{A}})=\textbf{diag}(\widehat{{\mathcal{A}}_{1,i}})+j\textbf{diag}(\widehat{{\mathcal{A}}_{j,k}})\in\mathbb{H}_{c}^{n_{3}n_{1}\times n_{3}n_{2}}.$ Its conjugate transpose is $(\textbf{diag}(\hat{\mathcal{A}}))^{\ast}=(\textbf{diag}(\widehat{{\mathcal{A}}_{1,i}}))^{\ast}+j(\textbf{diag}(\widehat{{\mathcal{A}}_{j,k}}))^{\ast}.$ Combing with equality (5) and Lemma 2.12, we have $\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$blacksquare
Corollary 2.14
$\mathcal{A}\in\mathbb{H}_{c}^{n\times n\times m}$ is Hermitian (resp. unitary) if and only if $\textbf{diag}(\widehat{\mathcal{A}})$ is Hermitian (resp. unitary).
Proof.  By Definition 2.8, it is clear that $\textbf{diag}(\widehat{\mathcal{I}})=I_{nm}.$ Then, by Theorem 2.3 and Proposition 2.13, the result follows. $\blacksquare$
Proposition 2.15
Given $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$, we have $(\mathcal{A}\ast_{Ht}\mathcal{B})^{\ast}=\mathcal{B}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}.$
Proof.  From Theorem 2.3 and Proposition 2.13, we only need to show
$(\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}}))^{\ast}=(\textbf{diag}(\widehat{\mathcal{B}}))^{\ast}(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$
In fact, this equality holds for matrices $\textbf{diag}(\widehat{\mathcal{A}})$ and $\textbf{diag}(\widehat{\mathcal{B}})$ over $\mathbb{H}_{c}$. $\blacksquare$
3 The Ht-SVD of third-order tensors over $\mathbb{H}_{c}$
The SVD of a matrix over $\mathbb{H}_{c}$ has been discussed in some papers (see, e.g., [13]). Let $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times n}$. If
$A_{(c),1}=U_{1}\Sigma_{1}V_{1}^{\ast}$ and $A_{(c),2}=U_{2}\Sigma_{2}V_{2}^{\ast},$
then the SVD of $A$ is
$A=U\Sigma V^{\ast},$ (6)
where $U=U_{1}e_{1}+U_{2}e_{2}$ and $V=V_{1}e_{1}+V_{2}e_{2}$ are unitary matrices, $\Sigma=\Sigma_{1}e_{1}+\Sigma_{2}e_{2}$ is diagonal but not a real matrix in general.
Recall that a tensor is called $f$-diagonal if each frontal slice is diagonal ([6]). Next, we propose the Ht-SVD of a reduced biquaternion tensor.
Theorem 3.1
(Ht-SVD) Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then there exist unitary tensors $\mathcal{U}\in\mathbb{H}_{c}^{n_{1}\times n_{1}\times n_{3}}$ and $\mathcal{V}\in\mathbb{H}_{c}^{n_{2}\times n_{2}\times n_{3}}$ such that
$\mathcal{A}=\mathcal{U}\ast_{Ht}\mathcal{S}\ast_{Ht}\mathcal{V}^{\ast},$
where $\mathcal{S}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ is $f$-diagonal. Such decomposition is called $Ht$-SVD of $\mathcal{A}$.
Proof.  We first transform $\mathcal{A}$ into $\widehat{\mathcal{A}}$ by using DFT along the third mode, and then obtain the block diagonal matrix $\textbf{diag}(\widehat{\mathcal{A}})$ by (1). In addition, we consider the SVD of each $\widehat{A}^{(s)}$ in $\textbf{diag}(\widehat{\mathcal{A}})$ with formula (6). Suppose $\widehat{A}^{(s)}=\widehat{U}^{(s)}\widehat{\Sigma}^{(s)}(\widehat{V}^{(s)})^{\ast},s=1,2,\cdots,n_{3}.$ Then we have
$\textbf{diag}(\widehat{\mathcal{A}})=\begin{bmatrix}\widehat{U}^{(1)}&&&\\ &\widehat{U}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{U}^{(n_{3})}\\ \end{bmatrix}\begin{bmatrix}\widehat{\Sigma}^{(1)}&&&\\ &\widehat{\Sigma}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{\Sigma}^{(n_{3})}\\ \end{bmatrix}\begin{bmatrix}(\widehat{V}^{(1)})^{\ast}&&&\\ &(\widehat{V}^{(2)})^{\ast}&&\\ &&\ddots&\\ &&&(\widehat{V}^{(n_{3})})^{\ast}\\ \end{bmatrix}.$
Setting
$\textbf{diag}(\widehat{\mathcal{U}})=\begin{bmatrix}\widehat{U}^{(1)}&&&\\ &\widehat{U}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{U}^{(n_{3})}\\ \end{bmatrix},\quad\textbf{diag}(\widehat{\mathcal{V}})=\begin{bmatrix}\widehat{V}^{(1)}&&&\\ &\widehat{V}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{V}^{(n_{3})}\\ \end{bmatrix},$
and
$\textbf{diag}(\hat{\mathcal{S}})=\begin{bmatrix}\widehat{\Sigma}^{(1)}&&&\\ &\widehat{\Sigma}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{\Sigma}^{(n_{3})}\\ \end{bmatrix},$
we have
$\textbf{diag}(\widehat{\mathcal{A}})=\textbf{diag}(\widehat{\mathcal{U}})\textbf{diag}(\widehat{\mathcal{S}})(\textbf{diag}(\widehat{\mathcal{V}}))^{\ast},$ (7)
in which $\textbf{diag}(\widehat{\mathcal{U}})$ and $\textbf{diag}(\widehat{\mathcal{V}})$ are unitary matrices, and $\textbf{diag}(\widehat{\mathcal{S}})$ is a block diagonal matrix with elements being diagonal matrices in general. Next, we construct reduced biquaternion tensors $\mathcal{U},\mathcal{V}$ and $\mathcal{S}$ as follows:
$\mathcal{U}\doteq\textbf{Fold}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})\textbf{diag}(\widehat{\mathcal{U}})(e\otimes I_{n_{1}})\right),$
$\mathcal{V}\doteq\textbf{Fold}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{2}})\textbf{diag}(\widehat{\mathcal{V}})(e\otimes I_{n_{2}})\right),$
$\mathcal{S}\doteq\textbf{Fold}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})\textbf{diag}(\widehat{\mathcal{S}})(e\otimes I_{n_{2}})\right).$
Observing that $\textbf{Vec}(\widehat{\mathcal{U}})=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{U}),\textbf{Vec}(\widehat{\mathcal{V}})=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{2}})\textbf{Vec}(\mathcal{V}),\textbf{Vec}(\widehat{\mathcal{S}})=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{S}).$ That is, $\widehat{\mathcal{U}},\widehat{\mathcal{V}},\widehat{\mathcal{S}}$ are the DFT of tensors $\mathcal{U},\mathcal{V},\mathcal{S}$, respectively. Thus by applying Theorem 2.3 and Proposition 2.13 , we infer from the equation (7) that
$\mathcal{A}=\mathcal{U}\ast_{Ht}\mathcal{S}\ast_{Ht}\mathcal{V}^{\ast}.$
Obviously, $\mathcal{S}$ is $f$-diagonal. And according to Corollary 2.14, we know that $\mathcal{U}$, $\mathcal{V}$ are our required unitary tensors. $\blacksquare$
4 The MP inverses of reduced biquaternion tensors
We refer the readers to [1] for Moore-Penrose (MP) inverses. In this section, as an application of Ht-SVD of a RB tensor, we define the MP inverse for a RB tensor and find a formula for the MP inverse by using Theorem 3.1. Moreover, we explore some theoretical properties of the MP inverse of RB tensors.
For a given matrix $A$, its MP inverse is the unique matrix $X$ (denoted as $A^{\dagger}$) satisfying
$AXA=A,\ XAX=X,\ (AX)^{\ast}=AX,\ (XA)^{\ast}=XA.$
According to the properties of reduced biquaternion matrices, we can easily derive the MP inverse of a reduced biquaternion matrix as follows.
Theorem 4.1
Let $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times n},$ where $A_{(c),1},\ A_{(c),2}\in\mathbb{C}^{m\times n}$. The MP inverse of $A$ is $A^{\dagger}=A_{(c),1}^{\dagger}e_{1}+A_{(c),2}^{\dagger}e_{2}\in\mathbb{H}_{c}^{n\times m}.$
Now, we introduce the definition of the MP inverse of a third-order Segre quaternion tensor.
Definition 4.2
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. If there exists $\mathcal{X}\in\mathbb{H}_{c}^{n_{2}\times n_{1}\times n_{3}}$ satisfying:
  • (i)
    $\mathcal{A}\ast_{Ht}\mathcal{X}\ast_{Ht}\mathcal{A}=\mathcal{A};$
  • (ii)
    $\mathcal{X}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{X};$
  • (iii)
    $(\mathcal{A}\ast_{Ht}\mathcal{X})^{\ast}=\mathcal{A}\ast_{Ht}\mathcal{X};$
  • (iv)
    $(\mathcal{X}\ast_{Ht}\mathcal{A})^{\ast}=\mathcal{X}\ast_{Ht}\mathcal{A},$
then $\mathcal{X}$ is said to be an MP inverse of $\mathcal{A}$, denoted by $\mathcal{A}^{\dagger}.$
For each $n_{1}\times n_{2}\times n_{3}$ reduced biquaternion tensor, the following theorem shows that it has a unique MP inverse.
Theorem 4.3
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and its Ht-SVD be $\mathcal{A}=\mathcal{U}\ast_{Ht}\mathcal{S}\ast_{Ht}\mathcal{V}^{\ast},$ where $\mathcal{U}\in\mathbb{H}_{c}^{n_{1}\times n_{1}\times n_{3}},\mathcal{V}\in\mathbb{H}_{c}^{n_{2}\times n_{2}\times n_{3}}$ are unitary and $\mathcal{S}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ is f-diagonal. Then $\mathcal{A}$ has a unique MP inverse
$\mathcal{A}^{\dagger}=\mathcal{V}\ast_{Ht}\mathcal{S}^{\dagger}\ast_{Ht}\mathcal{U}^{\ast},$ (8)
where $\mathcal{S}^{\dagger}=\textbf{Fold}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{2}})(\textbf{diag}(\widehat{\mathcal{S}}))^{\dagger}(e\otimes I_{n_{1}})\right)$.
Proof.  From Theorem 2.3, Proposition 2.5, Proposition 2.15 and Definition 4.1, it is not difficult to verify that $\textbf{diag}(\widehat{\mathcal{S}^{\dagger}})=(\textbf{diag}(\widehat{\mathcal{S}}))^{\dagger}$ and the expression of $\mathcal{A}^{\dagger}$ given in (8) satisfies all four conditions in Definition 4.2. Next, we need to prove the uniqueness of the MP inverse. Suppose that $\mathcal{X}$ and $\mathcal{Y}$ are two MP inverses of $\mathcal{A}$, respectively. Then
$\mathcal{X}$ $=\mathcal{X}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{X}\ast_{Ht}\mathcal{X}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}$
$=\mathcal{X}\ast_{Ht}\mathcal{X}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}\mathcal{Y}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}$
$=\mathcal{X}\ast_{Ht}\mathcal{Y}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{X}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Y}$
$=\mathcal{X}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Y}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Y}$
$=(\mathcal{X}\ast_{Ht}\mathcal{A})^{\ast}\ast_{Ht}(\mathcal{Y}\ast_{Ht}\mathcal{A})^{\ast}\ast_{Ht}\mathcal{Y}$
$=\mathcal{A}^{\ast}\ast_{Ht}\mathcal{Y}^{\ast}\ast_{Ht}\mathcal{Y}=\mathcal{Y}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Y}$
$=\mathcal{Y}.$
The proof is completed. $\blacksquare$
Next, we explore some properties of the MP inverse of a reduced biquaternion tensor $\mathcal{A}$.
Proposition 4.4
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then
  • (a)
    $(\mathcal{A}^{\dagger})^{\dagger}=\mathcal{A}$;
  • (b)
    $(\mathcal{A}^{\dagger})^{\ast}=(\mathcal{A}^{\ast})^{\dagger}$;
  • (c)
    $(\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast})^{\dagger}=(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}\mathcal{A}^{\dagger}$ and $(\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A})^{\dagger}=\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}^{\ast})^{\dagger}$;
  • (d)
    $(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})^{\dagger}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}$ ( $\mathcal{P}$ and $\mathcal{Q}$ are unitary reduced biquaternion tensors );
  • (e)
    If $\mathcal{A}=(a_{i_{1}i_{2}i_{3}})\in\mathbb{H}_{c}^{n_{1}\times n_{1}\times n_{3}}$ is f-diagonal, then $\mathcal{A}^{\dagger}$ is f-diagonal, and $(\mathcal{A}^{\dagger}_{i_{1}i_{2}i_{3}})=(a_{i_{1}i_{2}i_{3}})^{\dagger}$. In this case, if each diagonal entry in each frontal slice has an inverse, then $\mathcal{A}^{\dagger}=\mathcal{A}^{-1}$;
  • (f)
    $\mathcal{A}^{\dagger}=\mathcal{A}^{\ast}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast})^{\dagger}=(\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A})^{\dagger}\ast_{Ht}\mathcal{A}^{\ast}$;
  • (g)
    $\mathcal{A}=\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}(\mathcal{A}^{\ast})^{\dagger}=(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}$;
  • (h)
    $\mathcal{A}^{\ast}=\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}$;
  • (i)
    $\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}=(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}\mathcal{A}^{\ast}$ and $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}=\mathcal{A}^{\ast}\ast_{Ht}(\mathcal{A}^{\ast})^{\dagger}$;
  • (j)
    $\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}=(\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast})^{\dagger}$;
  • (k)
    $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}=(\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A})^{\dagger}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}=\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}\ast_{Ht}(\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A})^{\dagger}$;
  • (l)
    $\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}$ if $\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}$;
  • (m)
    $\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}$ and $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}$ are idempotent;
  • (n)
    If $\mathcal{A}$ is Hermitian and idempotent, then $\mathcal{A}^{\dagger}=\mathcal{A}$.
Proof.  We only prove (b), (c), (d), (l) and (n). In the analogous way, one can obtain the remaining properties.
  • (b)
    Using Propositions 2.5 and 2.15, one can obtain that $(\mathcal{A}^{\dagger})^{\ast}$ is the MP inverse of $\mathcal{A}^{\ast}$ by direct calculations.
  • (c)
    By (b), Proposition 2.5, Proposition 2.15 and the fact $(\mathcal{A}^{\ast})^{\ast}=\mathcal{A}$, one can get that $(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}\mathcal{A}^{\dagger}$ is the MP inverse of $\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}$ and $\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}^{\ast})^{\dagger}$ is the MP inverse of $\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}$.
  • (d)
    Suppose $\mathcal{P}\ast_{Ht}\mathcal{P}^{\ast}=\mathcal{P}^{\ast}\ast_{Ht}\mathcal{P}=\mathcal{I}$ and $\mathcal{Q}\ast_{Ht}\mathcal{Q}^{\ast}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{Q}=\mathcal{I}$. Then we have
    $\begin{array}[]{lll}&(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})\ast_{Ht}(\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast})\ast_{Ht}(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})\\ &=\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q}=\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q},\end{array}$
    $\begin{array}[]{lll}&(\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast})\ast_{Ht}(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})\ast_{Ht}(\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast})\\ &=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast},\end{array}$
    $\begin{array}[]{lll}&(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q}\ast_{Ht}\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast})^{\ast}\\ &=\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}=\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q}\ast_{Ht}\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast},\end{array}$
    and
    $(\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}\ast_{Ht}\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})^{\ast}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}\ast_{Ht}\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q}.$
    Hence, by Definition 4.2 we can derive that $(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})^{\dagger}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}.$
  • (l)
    It follows from (i) and (j).
  • (n)
    If $\mathcal{A}$ is Hermitian and idempotent, then $\mathcal{A}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}=\mathcal{A}$ and $(\mathcal{A}\ast_{Ht}\mathcal{A})^{\ast}=\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{A}\ast_{Ht}\mathcal{A}$. Thus by Definition 4.2 we have $\mathcal{A}^{\dagger}=\mathcal{A}$.
$\blacksquare$
In general, $(\mathcal{A}\ast_{Ht}\mathcal{B})^{\dagger}\neq\mathcal{B}^{\dagger}\ast_{Ht}\mathcal{A}^{\dagger}$. Under some conditions, we have the equality case given as follows.
Proposition 4.5
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$. Then $(\mathcal{A}\ast_{Ht}\mathcal{B})^{\dagger}=\mathcal{B}^{\dagger}\ast_{Ht}\mathcal{A}^{\dagger}$ if one of the following conditions satisfies.
(a) $\mathcal{A}=\mathcal{B}^{\dagger}$;   (b) $\mathcal{A}=\mathcal{B}^{\ast}$;  (c) $\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}=\mathcal{I}$;  (d) $\mathcal{B}\ast_{Ht}\mathcal{B}^{\ast}=\mathcal{I}$.
Denote $\mathcal{L}_{\mathcal{A}}=\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}\mathcal{A},\ \mathcal{R}_{\mathcal{A}}=\mathcal{I}-\mathcal{A}\ast_{Ht}{\mathcal{A}}^{\dagger}.$ $\mathcal{O}$ represents zero tensor with all the entries being zero. Obviously, $\mathcal{A}\ast_{Ht}\mathcal{O}=\mathcal{O}\ast_{Ht}\mathcal{A}=\mathcal{O}.$
Proposition 4.6
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then
  • (a)
    $\mathcal{A}\ast_{Ht}\mathcal{L}_{\mathcal{A}}=\mathcal{O}$ and $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{A}=\mathcal{O}$;
  • (b)
    $\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{A}^{\dagger}=\mathcal{O}$ and $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{R}_{\mathcal{A}}=\mathcal{O}$;
  • (c)
    $\mathcal{L}_{\mathcal{A}^{\ast}}=(\mathcal{R}_{\mathcal{A}})^{\ast}=\mathcal{R}_{\mathcal{A}}$ and $\mathcal{R}_{\mathcal{A}^{\ast}}=(\mathcal{L}_{\mathcal{A}})^{\ast}=\mathcal{L}_{\mathcal{A}}$;
  • (d)
    $\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{L}_{\mathcal{A}}=\mathcal{L}_{\mathcal{A}}$ and $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{R}_{\mathcal{A}}=\mathcal{R}_{\mathcal{A}}$;
  • (e)
    $(\mathcal{L}_{\mathcal{A}})^{\dagger}=\mathcal{L}_{\mathcal{A}}$ and $(\mathcal{R}_{\mathcal{A}})^{\dagger}=\mathcal{R}_{\mathcal{A}}$;
  • (f)
    $\mathcal{L}_{\mathcal{A}}=\mathcal{L}_{\mathcal{A}^{\ast}\ast_{Ht}{\mathcal{A}}}$ and $\mathcal{R}_{\mathcal{A}}=\mathcal{R}_{\mathcal{A}\ast_{Ht}{\mathcal{A}^{\ast}}}$.
Proof.  By straightforward calculations, one can prove that (a), (b) and (d) hold.
For (c), the first equality follows from
$\mathcal{L}_{\mathcal{A}^{\ast}}=\mathcal{I}-(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}{\mathcal{A}^{\ast}}=\mathcal{I}-(\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger})^{\ast}=(\mathcal{I}-\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger})^{\ast}=(\mathcal{R}_{\mathcal{A}})^{\ast}$
and
$\mathcal{L}_{\mathcal{A}^{\ast}}=\mathcal{I}-(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}{\mathcal{A}^{\ast}}=\mathcal{I}-(\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger})^{\ast}=\mathcal{I}-\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}=\mathcal{R}_{\mathcal{A}}.$
In an analogous manner, one can prove the second part.
For (e), it follows immediately from (c), (d) and Proposition 4.4.
For (f), $\mathcal{L}_{\mathcal{A}^{\ast}\ast_{Ht}{\mathcal{A}}}=\mathcal{I}-(\mathcal{A}^{\ast}\ast_{Ht}{\mathcal{A}})^{\dagger}\ast_{Ht}(\mathcal{A}^{\ast}\ast_{Ht}{\mathcal{A}})=\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}({\mathcal{A}}^{\dagger})^{\ast}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}{\mathcal{A}}=\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}{\mathcal{A}}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}{\mathcal{A}}=\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}{\mathcal{A}}=\mathcal{L}_{\mathcal{A}}.$ The second part can be proved in a similar way. $\blacksquare$
5 Reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$
We first discuss the consistency of RB tensor equation
$\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$ (9)
by using the MP inverse demonstrated in Section 4. The solvability conditions are found and a general expression of the solutions is provided when the solutions exist. Furthermore, we also derive solvability conditions and the general solutions involving Hermicity.
Theorem 5.1
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}.$ Then equation (9) has solutions if and only if $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}$.
Furthermore, the general solutions to equation (9) can be written as
$\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{Y},$ (10)
with $\mathcal{Y}$ be any reduced biquaternion tensor with compatible size.
Proof.  Assume that $\mathcal{C}$ is a solution to reduced biquaternion tensor equation (9). Then by Proposition 2.5 and Proposition 4.6, we have
$\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{C}=\mathcal{O}.$
Conversely, if $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}$, it is easy to see that $\mathcal{X}_{0}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}$ is a solution to equation (9), and so is $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{Y}$.
Moreover, each solution $\mathcal{X}$ of equation (9) has the form of the equation (10) since $\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}\mathcal{A})\ast_{Ht}\mathcal{X}.$blacksquare
Theorem 5.2
Let $\mathcal{A},\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then equation (9) has a Hermitian solution if and only if $\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast}=\mathcal{B}\ast_{Ht}\mathcal{A}^{\ast}$ and $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}.$
Moreover, if (9) is solvable, then the general Hermitian solutions to the Segre quaternion tensor equation (9) can be written as
$\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B})^{\ast}-\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast})\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{U}\ast_{Ht}\mathcal{L}_{\mathcal{A}},$ (11)
where $\mathcal{U}=\mathcal{U}^{\ast}$ is an arbitrary tensor with appropriate size.
Proof.  If equation (9) has a Hermitian solution $\mathcal{C}$, then by Propositions 2.5, 2.15 and Definition 4.2, we obtain
$\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast}=\mathcal{A}\ast_{Ht}\mathcal{C}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{A}\ast_{Ht}\mathcal{C}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{B}\ast_{Ht}\mathcal{A}^{\ast},$
and $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{C}=\mathcal{O}.$
On the other hand, assume that $\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast}=\mathcal{B}\ast_{Ht}\mathcal{A}^{\ast}$ and $R_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}$ hold. Then by applying Propositions 2.5 and 2.15, we get that
$\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B})^{\ast}-\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast})\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{U}\ast_{Ht}\mathcal{L}_{\mathcal{A}}$
is a Hermitian solution to Segre quaternion tensor equation (9).
Next we show that each solution of Segre quaternion tensor equation (9) should be in the form of (11). Assume $\mathcal{X}_{0}$ is an arbitrary Hermitian solution to tensor equation (9). Setting $\mathcal{U}=\mathcal{X}_{0}$ yields
$\mathcal{X}=$ $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B})^{\ast}-\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast})\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{X}_{0}\ast_{Ht}\mathcal{L}_{\mathcal{A}}$
$=$ $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B})^{\ast}-\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast})\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}$
$+(\mathcal{X}_{0}-\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{X}_{0})\ast_{Ht}(\mathcal{I}-\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A})$
$=$ $(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B})^{\ast}-\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast})\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}$
$+\mathcal{X}_{0}-\mathcal{X}_{0}\ast_{Ht}(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A})+\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}$
$=$ $(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{X}_{0})^{\ast}+\mathcal{X}_{0}-\mathcal{X}_{0}\ast_{Ht}(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A})^{\ast}$
$=$ $\mathcal{X}_{0},$
which implies that each Hermitian solution to Segre quaternion tensor equation (9) can be represented as (11). Thus, (11) is the general expression of Hermitian solutions to equation (9). $\blacksquare$
Next, we deal with the least-squares problem of $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$. It is well known that Segre quaternion matrices have the following properties. The equivalent complex matrix of reduced biquaternion matrix $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times n}$ ([13]) is
$\widetilde{A}_{(c)}=\begin{bmatrix}A_{(c),1}&0\\ 0&A_{(c),2}\\ \end{bmatrix}.$ (12)
Given $A,B\in\mathbb{H}_{c}^{m\times n}$ and $C\in\mathbb{H}_{c}^{n\times l}$, we have $(\widetilde{A+B})_{(c)}=\widetilde{A}_{(c)}+\widetilde{B}_{(c)}$ and $(\widetilde{AC})_{(c)}=\widetilde{A}_{(c)}\widetilde{C}_{(c)}$. For $A=(a_{i_{1}i_{2}})\in\mathbb{H}_{c}^{m\times n}$, we define the Frobenius norm of $A$ as $\|A\|_{F}=\sqrt{\sum_{i_{1}=1}^{m}\sum_{i_{2}=1}^{n}|a_{i_{1}i_{2}}|^{2}}$, and $\infty$-norm as $\|A\|_{\infty}=\mathop{\max}_{1\leq i\leq m}\sum_{j=1}^{n}|a_{ij}|$. Moreover, it can be verified that
$\|A\|_{F}=\frac{1}{\sqrt{2}}\sqrt{\|A_{(c),1}\|^{2}+\|A_{(c),2}\|^{2}}=\frac{1}{\sqrt{2}}\|\widetilde{A}_{(c)}\|_{F}.$ (13)
$\|A\|_{\infty}=\mathop{\max}_{1\leq i\leq n_{1}n_{3}}\sum_{j=1}^{{n_{2}}}(|c_{ij}|+|d_{ij}|),$
where $A_{(c),1}=(c_{ij})_{m\times n},$ $A_{(c),2}=(d_{ij})_{m\times n}.$
For a tensor $\mathcal{A}=\mathcal{A}_{1,i}+j\mathcal{A}_{j,k}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ with $\mathcal{A}_{1,i},\mathcal{A}_{j,k}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$, we define its Frobenius norm as
$\|\mathcal{A}\|_{F}^{2}=\|\textbf{Vec}(\mathcal{A})\|_{F}^{2}=\|\textbf{Vec}(\mathcal{A}_{1,i})\|_{F}^{2}+\|\textbf{Vec}(\mathcal{A}_{j,k})\|_{F}^{2}.$ (14)
$\|\mathcal{A}\|_{\infty}=\mathop{\max}_{1\leq i\leq n_{1}n_{3}}\sum_{j=1}^{{n_{2}}}(|c_{ij}|+|d_{ij}|),$ (15)
where $\textbf{Vec}(\mathcal{A}_{1,i})=(c_{ij})_{n_{1}n_{3}\times{n_{2}}},$ $\textbf{Vec}(\mathcal{A}_{j,k})=(d_{ij})_{n_{1}n_{3}\times{n_{2}}}.$
Theorem 5.3
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}.$ Then $\|\mathcal{A}\|_{F}=\frac{1}{\sqrt{n_{3}}}\|\textbf{diag}(\hat{\mathcal{A}})\|_{F}$.
Proof.  By applying identities (3), (14) and $F_{n_{3}}j=jF_{n_{3}}$ as well as the invariant property of Frobenius norm under a unitary transformation for complex matrices, we have
$\|\textbf{diag}(\hat{\mathcal{A}})\|_{F}^{2}$ $=\|\textbf{Vec}(\hat{\mathcal{A}})\|_{F}^{2}$
$=\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{1,i})+\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})j\textbf{Vec}(\mathcal{A}_{j,k})\|_{F}^{2}$
$=\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{1,i})\|_{F}^{2}+\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{j,k})\|_{F}^{2}$
$=n_{3}\|\textbf{Vec}(\mathcal{A}_{1,i})\|_{F}^{2}+n_{3}\|\textbf{Vec}(\mathcal{A}_{j,k})\|_{F}^{2}$
$=n_{3}\|\textbf{Vec}(\mathcal{A})\|_{F}^{2}$
$=n_{3}\|\mathcal{A}\|_{F}^{2}.$
Therefore, $\|\mathcal{A}\|_{F}=\frac{1}{\sqrt{n_{3}}}\|\textbf{diag}(\hat{\mathcal{A}})\|_{F}$. $\blacksquare$
Lemma 5.4
([1]) The solutions of least-squares problem of complex equation $AX=B$ are
$X=A^{\dagger}B+(I-A^{\dagger}A)Y,$
with $Y$ be any matrix with appropriate size. Furthermore, $A^{\dagger}B$ is the minimum norm least-squares solution.
Lemma 5.5
Suppose $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times n},\ B=B_{(c),1}e_{1}+B_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times l},$ Denote $\dot{B}:=\begin{bmatrix}B_{(c),1}\\ B_{(c),2}\\ \end{bmatrix}.$ Then reduced biquaternion least-squares problem $AX=B$ has a solution $X=X_{(c),1}e_{1}+X_{(c),2}e_{2}\in\mathbb{H}_{c}^{n\times l}$ if and only if the complex least-squares problem $\widetilde{A}_{(c)}\dot{X}=\dot{B}$ has a solution $\dot{X}=\begin{bmatrix}X_{(c),1}\\ X_{(c),2}\\ \end{bmatrix}$, where $\widetilde{A}_{(c)}$ is given by (12).
In addition, if $\dot{X}$ is the minimum norm solution to the complex least-squares problem $\widetilde{A}_{(c)}\dot{X}=\dot{B}$, then $X=\begin{bmatrix}I_{n},&0\\ \end{bmatrix}\dot{X}e_{1}+\begin{bmatrix}0,&I_{n}\\ \end{bmatrix}\dot{X}e_{2}$ is the minimum norm solution to the least-squares problem $AX=B$ over $\mathbb{H}_{c}$.
Proof.  By equation (13), we have
$\|AX-B\|_{F}=\frac{1}{\sqrt{2}}\|\widetilde{A}_{(c)}\widetilde{X}_{(c)}-\widetilde{B}_{(c)}\|_{F},$
where $\widetilde{A}_{(c)},\widetilde{B}_{(c)}$ and $\widetilde{X}_{(c)}$ are defined by (12). Then using the fact
$\|\widetilde{A}_{(c)}\widetilde{X}_{(c)}-\widetilde{B}_{(c)}\|_{F}=\|\widetilde{A}_{(c)}\begin{bmatrix}X_{(c),1}\\ X_{(c),2}\\ \end{bmatrix}-\begin{bmatrix}B_{(c),1}\\ B_{(c),2}\\ \end{bmatrix}\|_{F},$
we can see that $\|AX-B\|_{F}$ and $\|\widetilde{A}_{(c)}\dot{X}-\dot{B}\|_{F}$ arrive the minima at the same time. With Lemma 5.4,
$\dot{X}=(\widetilde{A}_{(c)})^{\dagger}\dot{B}+(I-(\widetilde{A}_{(c)})^{\dagger}\widetilde{A}_{(c)})Y$
is the least-squares solution to $\widetilde{A}_{(c)}\dot{X}=\dot{B}$, in which $Y$ is arbitrary, and $X=\begin{bmatrix}I_{n},&0\\ \end{bmatrix}\dot{X}e_{1}+\begin{bmatrix}0,&I_{n}\\ \end{bmatrix}\dot{X}e_{2}$ are the least-squares solutions to $AX=B.$blacksquare
Theorem 5.6
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}$ with their DFTs $\hat{\mathcal{A}}$ and $-^\hat{\mathcal{B}}$. Denote $-^\textbf{A}\doteq\textbf{diag}(\hat{\mathcal{A}}),\textbf{B}\doteq\textbf{Vec}(\hat{\mathcal{B}}).$ Then $\mathcal{X}\in\mathbb{H}^{n_{2}\times n_{4}\times n_{3}}$ is a solution to the tensor least-squares problem $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$ if and only if $\textbf{X}=\textbf{Vec}(\hat{\mathcal{X}})$ is a solution to the matrix least-squares problem $\textbf{A}\textbf{X}=\textbf{B}.$
Moreover, if X is the minimum norm least-squares solution to $\textbf{A}\textbf{X}=\textbf{B}$, then
$=\textbf{Fold}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{2}})\textbf{X}\right)$ X )
is the minimum norm least-squares solution to $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}.$
Proof.  By Theorem 2.3 and Theorem 5.3, we have
$\|\mathcal{A}\ast_{Ht}\mathcal{X}-\mathcal{B}\|_{F}=\frac{1}{\sqrt{n_{3}}}\|\textbf{diag}(\hat{\mathcal{A}})\textbf{diag}(\hat{\mathcal{X}})-\textbf{diag}(\hat{\mathcal{B}})\|_{F}.$
[…]
6 Applications in color video compression
We use a reduced biquaternion tensor $\mathcal{\tilde{T}}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ to express a color video, i.e.,
$\mathcal{\tilde{T}}_{C}(m,n,t)=\mathcal{\tilde{T}}_{R}(m,n,t)i+\mathcal{\tilde{T}}_{G}(m,n,t)j+\mathcal{\tilde{T}}_{B}(m,n,t)k,$
where the red $\mathcal{\tilde{T}}_{R}(m,n,t)$, green $\mathcal{\tilde{T}}_{G}(m,n,t)$ and blue $\mathcal{\tilde{T}}_{B}(m,n,t)$ pixel values are sets of the t-th color image of a color video. By Theorem 3.1, we can get the Ht-SVD of the $\mathcal{\tilde{T}}$.
For comparison, we apply the Ht-SVD method to compress the same color video “DO01_013” as the one in [15]. While in [15], it was compressed by Qt-SVD. Figure 1 shows the PSNRs of the compressed video when taking the rank-k (k=10, 20, 50) approximations to “DO01_013” for the 1st, the 20th and the 50th frames, respectively, where PSNR is defined as
$PSNR\doteq 10log_{10}\left(\frac{3n_{1}n_{2}\|\mathcal{\tilde{T}}\|_{\infty}^{2}}{\|\mathcal{\tilde{T}}-\mathcal{T}\|_{F}^{2}}\right),$
with $\mathcal{T},\mathcal{\tilde{T}}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{T}$ is the approximated tensor of $\mathcal{\tilde{T}}$.
To implement the Ht-SVD method, one needs to calculate $2n_{3}$ SVDs of complex matrices with size of ${n_{1}\times n_{2}}$. While by the Qt-SVD method, one needs to calculate $n_{3}$ QSVDs of quaternion matrices with size of ${n_{1}\times n_{2}}$. From Fig. 1, we see that the performance is comparable to Fig. 2 in [15]. But Table 1 shows that the using of the Ht-SVD is much faster than Qt-SVD in [15].
Figure 1: PSNRs of the rank-k (k=10, 20, 50) approximations to DO01_013 by Ht-SVD .
frame 10 20 30 40
Qt-SVD 20.861257s 46.294602s 89.366687s 116.283031s
Ht-SVD 2.715995s 6.818766s 9.597188s 13.1181154s
Table 1: The running time of Ht-SVD and Qt-SVD for the different number of frames