### v2

Abstract
In this paper, we introduce the applications of third-order reduced biquaternion tensors in color video processing. We first develop an algorithm for computing the singular value decomposition (SVD) of a third-order reduced biquaternion tensor via a new Ht-product. As theoretical applications, we define the Moore-Penrose inverse of a third-order reduced biquaternion tensor and consider its characterizations via its SVD. Using Moore-Penrose inverses, we mainly discuss the general (or Hermitian) solutions to reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$ as well as its least-squares solutions. Finally, we develop two algorithms and apply them in color video compression and deblurring, both of which perform better than the compared algorithms.
Key words: Reduced biquaternion tensor; Ht-product; Singular value decomposition; Moore-Penrose inverse; Color video processing.
AMS 2010 Subject Classification: 15A09, 15A18, 15A23, 15A24, 15B33, 39B42.
1 Introduction
In 1843, Hamilton extended the complex number field $\mathbb{C}$ to quaternions which is a four dimensional division algebra over the real number field $\mathbb{R}$. It is well-known that Hamilton quaternions have successful applications in signal and color image processing (e.g., [4]). To keep the commutativity of usual number systems, in 1892, Segre ([15]) discovered the following extension:
$\mathbb{H}_{c}=\{a_{1}+a_{2}\mathbf{i}+a_{3}\mathbf{j}+a_{4}\mathbf{k}\ |\ \mathbf{i}^{2}=-1,\ \mathbf{j}^{2}=1,\ \mathbf{ij}=\mathbf{ji}=\mathbf{k},\ a_{1},\dots,a_{4}\in\mathbb{R}\}.$
Clearly, $\mathbf{k}^{2}=-1,\ \mathbf{ki}=\mathbf{ik}=-\mathbf{j},\ \mathbf{jk}=\mathbf{kj}=\mathbf{i}$. $\mathbb{H}_{c}$ is a four dimensional commutative algebra over $\mathbb{R}$ and contains zero divisors, and it is often called Segre quaternions ([12]), commutative quaternions ([8, 10, 11]) or reduced biquaternions ([11, 14]). In this paper, we call it “reduced biquaternions” ( or RB for short).
Reduced biquaternions $\mathbb{H}_{c}$ has numerous applications in many areas. In signal and color image processing, Pei et al. ([10, 11]) showed that the operations of the discrete reduced biquaternion Fourier transforms and their corresponding convolutions and correlations were much simpler than the existing implementations of the Hamilton quaternions, and many questions could be performed simultaneously by $\mathbb{H}_{c}$. They used reduced biquaternion matrices to represent the color images, and then compressed the images by the SVDs of reduced biquaternion matrices (RBSVDs). For comparison, their proposed algorithm (RBSVDs) was faster and has better reconstruction quality than the compared algorithm of computing the SVD of a quaternion matrix. By adopting the special structure of a real representation of a reduced biquaternion matrix, Ding et al. [3] investigated the least-squares special minimal norm solutions to RB matrix equation $AX=B$ and applied the least-squares minimal norm RB solution to the color image restoration.
A tensor is an extension of the concept of a matrix, which provides a more general and flexible approach to handle and represent complex multidimensional data structures. Tensors and tensor decompositions ([2, 5, 6, 7, 9, 13, 16]) have been applied in many areas such as signal processing, computer science, data mining, neuroscience, etc. Studying tensor rank decomposition may go back to the 1960s: the CANDECOMP/PARAFAC (CP) decomposition ([2]) and the Tucker decomposition ([16]). In [5], Kilmer and Martin defined a closed multiplication operation (called t-product) for two real tensors, and considered some important theoretical properties and practical tensor approximation problems. Qin et al. [13] defined a tensor product for third-order quaternion tensors and discussed the Qt-SVD of a third-order quaternion tensor with applications in color video. We note that a RB tensor also has good capability in representing color video as it does not lose the correlation between the RGB channels. Thus this paper will explore some fundamental theories of RB tensors and their applications, such as utilizing the proposed theories of SVD, Moore-Penrose inverses, the minimal norm least-squares solution to compress or deblur color videos.
This paper is organized as follows. We first introduce some operations over reduced biquaternion tensors, and then we define the Ht-product and explore some important properties of third-order RB tensors in Section 2. In Section 3, we investigate the Ht-SVD of a third-order RB tensor and develop an algorithm to compute it. Two theoretical applications of Ht-SVD are given in Sections 4 and 5. That is, in Section 4, the Moore-Penrose inverse of a third-order RB tensor is defined and some properties including algebraic expression derived by the Ht-SVD are discussed. In Section 5, we mainly consider the general/Hertmian/least-squares solutions/minimal norm least-squares solution to reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$. Finally, in Section 6, we apply our results to color video compression and deblurring. The experimental data show that our methods have excellent performances comparing with other methods.
2 The Ht-product of third-order tensors over $\mathbb{H}_{c}$
Throughout this paper, Euler script letters $(\mathcal{A,B,\cdots})$ are used to refer to tensors, while the capital letters $(A,B,\cdots)$ represent matrices. The notations $\mathbb{R}^{n_{1}\times n_{2}}$ and $\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ represent the set of all the matrices of dimension $n_{1}\times n_{2}$ and all the third-order tensors of dimension $n_{1}\times n_{2}\times n_{3}$ over algebra $\mathbb{R}$, respectively. For a third-order tensor $\mathcal{A}$, we denote the $i$th horizontal, the $j$th lateral and the $k$th frontal slice by $\mathcal{A}(i,:,:),\mathcal{A}(:,j,:)$ and $\mathcal{A}(:,:,k)$ respectively. For simplicity, let $A^{(k)}$ represent the $k$th frontal slice $\mathcal{A}(:,:,k)$.
Recall that any element $q=q_{0}+q_{1}\mathbf{i}+q_{2}\mathbf{j}+q_{3}\mathbf{k}\in\mathbb{H}_{c}$ can be expressed as $q=q_{a}+\mathbf{j}q_{b}=q_{(c),1}e_{1}+q_{(c),2}e_{2}$, where $q_{a}=q_{0}+q_{1}\mathbf{i},q_{b}=q_{2}+q_{3}\mathbf{i},$ $q_{(c),1}=q_{a}+q_{b},\ q_{(c),2}=q_{a}-q_{b}$, and $e_{1}=(1+\mathbf{j})/2,e_{2}=(1-\mathbf{j})/2$. Clearly, $e_{1}$ and $e_{2}$ are two special reduced biquaternions satisfying $e_{1}=e_{1}^{2}=\cdots=e_{1}^{n},e_{2}=e_{2}^{2}=\cdots=e_{2}^{n}$ and $e_{1}e_{2}=0$ ([11]). The modulus of $q$ is $|q|=\sqrt{q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}}$.
For any matrix $A\in\mathbb{H}_{c}^{m\times n}$, we can write it in two forms, that is, $A=A_{0}+A_{1}\mathbf{i}+A_{2}\mathbf{j}+A_{3}\mathbf{k}=B_{1}+\mathbf{j}B_{2}$ with $A_{0},\dots,A_{3}\in\mathbb{R}^{m\times n},B_{1}=A_{0}+A_{1}\mathbf{i},B_{2}=A_{2}+A_{3}\mathbf{i}\in\mathbb{C}^{m\times n}$. Then its conjugate transpose is given by $A^{\ast}=A_{0}^{T}-A_{1}^{T}\mathbf{i}+A_{2}^{T}\mathbf{j}-A_{3}^{T}\mathbf{k}=B_{1}^{\ast}+\mathbf{j}B_{2}^{\ast}$, where $B_{1}^{\ast},B_{2}^{\ast}$ denote the conjugate transpose of complex matrices $B_{1},B_{2}$. Moreover, $A\in\mathbb{H}^{n\times n}_{c}$ is said to be Hermitian (resp. unitary) if and only if $A=A^{\ast}$ (resp. $AA^{\ast}=A^{\ast}A=I_{n}$). Furthermore, if we rewrite $A\in\mathbb{H}_{c}^{m\times n}$ as $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}$, where $A_{(c),1}=B_{1}+B_{2}$ and $A_{(c),2}=B_{1}-B_{2}$ are complex matrices, then the conjugate transpose of $A$ is $A^{\ast}=A_{(c),1}^{\ast}e_{1}+A_{(c),2}^{\ast}e_{2}$. Thus for any $B=B_{(c),1}e_{1}+B_{(c),2}e_{2}\in\mathbb{H}_{c}^{n\times l}$, $AB=A_{(c),1}B_{(c),1}e_{1}+A_{(c),2}B_{(c),2}e_{2}$, and we can use it to prove that $(AB)^{\ast}=B^{\ast}A^{\ast}$. For $A=(a_{i_{1}i_{2}})\in\mathbb{H}_{c}^{m\times n}$, we define the Frobenius norm of $A$ as $\|A\|_{F}=\sqrt{\sum_{i_{1}=1}^{m}\sum_{i_{2}=1}^{n}|a_{i_{1}i_{2}}|^{2}}$. Moreover, it can be verified that $\|A\|_{F}^{2}=\mbox{tr}(\Re(A^{\ast}A)),$ where $\Re(\cdot)$ denotes the real part of a reduced biquaternion matrix.
The block circulant matrix $\textbf{circ}(\mathcal{A})$ generated by a third-order tensor $\mathcal{A}$’s frontal slices $A^{(1)},$ $A^{(2)},\cdots,A^{(n_{3})}$ is given as
$\textbf{circ}(\mathcal{A})=\begin{bmatrix}A^{(1)}&A^{(n_{3})}&\cdots&A^{(2)}\\ A^{(2)}&A^{(1)}&\cdots&A^{(3)}\\ \vdots&\vdots&\ddots&\vdots\\ A^{(n_{3})}&A^{(n_{3}-1)}&\cdots&A^{(1)}\\ \end{bmatrix}.$
We will use Vec operator to convert a tensor $\mathcal{A}$ into a block matrix $\textbf{Vec}(\mathcal{A})$:
$\textbf{Vec}(\mathcal{A})=\begin{bmatrix}A^{(1)}\\ A^{(2)}\\ \vdots\\ A^{(n_{3})}\\ \end{bmatrix},$
and use the inverse operation Fold to take the block matrix $\textbf{Vec}(\mathcal{A})$ back to the tensor $\mathcal{A}$:
$\textbf{Fold}(\textbf{Vec}(\mathcal{A}))=\mathcal{A}.$
Moreover, we use the frontal slices of $\mathcal{A}$ to define a block diagonal matrix as following:
$\textbf{diag}({\mathcal{A}})\doteq\begin{bmatrix}{A}^{(1)}&&&\\ &{A}^{(2)}&&\\ &&\ddots&\\ &&&{A}^{(n_{3})}\\ \end{bmatrix}.$ (1)
Note that by using the normalized Discrete Fourier transformation (DFT) matrix, a complex circulant matrix can be diagonalized ([5], [13]). Suppose $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ is a third-order complex tensor, then
$(F_{n_{3}}\otimes I_{n_{1}})\textbf{circ}(\mathcal{A})(F_{n_{3}}^{\ast}\otimes I_{n_{2}})=\begin{bmatrix}\widehat{A}^{(1)}&&&\\ &\widehat{A}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{A}^{(n_{3})}\\ \end{bmatrix}=\textbf{diag}(\widehat{\mathcal{A}})\in\mathbb{C}^{n_{3}n_{1}\times n_{3}n_{2}},$ (2)
where $\otimes$ denotes the Kronecker product, $\widehat{A}^{(i)}$ ($i=1,\cdots,n_{3}$) are the frontal slices of the tensor $\widehat{\mathcal{A}}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ which is the result of DFT of $\mathcal{A}$ along the third mode, $F_{n_{3}}$ is an $n_{3}\times n_{3}$ unitary matrix defined by $F_{n_{3}}(s,t)=\frac{1}{\sqrt{n_{3}}}\omega^{(s-1)(t-1)},\ \omega=exp(-\frac{2\pi\mathbf{i}}{n_{3}})$.
Definition 2.1
For the tensor $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$, the DFT $\widehat{\mathcal{A}}$ of $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ along the third mode is defined as
$\widehat{\mathcal{A}}(s,t,:)=\sqrt{n_{3}}F_{n_{3}}\mathcal{A}(s,t,:).$
As we can see that after the DFT, the resulting vector $\widehat{\mathcal{A}}(s,t,:)$ is still a vector in $\mathbb{H}^{n_{3}}_{c}$. We can verify the relation between $\mathcal{A}$ and $\widehat{\mathcal{A}}$ as follows.
$\textbf{Vec}(\widehat{\mathcal{A}})=\begin{bmatrix}\widehat{A}^{(1)}\\ \widehat{A}^{(2)}\\ \vdots\\ \widehat{A}^{(n_{3})}\\ \end{bmatrix}=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}).$ (3)
Now we extend $t$-product of two third-order real tensors defined in [5] to Reduced biquaternions.
Definition 2.2
(Ht-product) Let $\mathcal{A}=\mathcal{A}_{1,\mathbf{i}}+\mathbf{j}\mathcal{A}_{\mathbf{j},\mathbf{k}}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$. Then $Ht$-product of $\mathcal{A}$ and $\mathcal{B}$ is defined as
$\mathcal{A}\ast_{Ht}\mathcal{B}=\textbf{Fold}((\textbf{circ}(\mathcal{A}_{1,\mathbf{i}})+\mathbf{j}\textbf{circ}(\mathcal{A}_{\mathbf{j},\mathbf{k}}))\textbf{Vec}(\mathcal{B}))\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}.$
The addition and scalar multiplication rules of tensors over $\mathbb{H}_{c}$ are defined in a usual way.
Theorem 2.3
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}},\ \mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$ and $\mathcal{C}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}$ with the corresponding DFTs $\widehat{\mathcal{A}},\ \widehat{\mathcal{B}}$ and $\widehat{\mathcal{C}}$. Then $\mathcal{A}\ast_{Ht}\mathcal{B}=\mathcal{C}$ if and only if $\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}})=\textbf{diag}(\widehat{\mathcal{C}}).$
Proof.  Since $\widehat{\mathcal{A}}$ is the DFT of $\mathcal{A}=\mathcal{A}_{1,\mathbf{i}}+\mathbf{j}\mathcal{A}_{\mathbf{j},\mathbf{k}}$ along the third mode, it follows from (3) and $F_{n_{3}}\mathbf{j}=\mathbf{j}F_{n_{3}}$ that
$\textbf{Vec}(\widehat{\mathcal{A}})$ $=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})(\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})+\mathbf{j}\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}}))$
$=\textbf{Vec}(\widehat{\mathcal{A}}_{1,\mathbf{i}})+\mathbf{j}\textbf{Vec}(\widehat{\mathcal{A}}_{\mathbf{j},\mathbf{k}}),$
which implies
$\textbf{diag}(\widehat{\mathcal{A}})=\textbf{diag}(\widehat{\mathcal{A}}_{1,\mathbf{i}})+\mathbf{j}\textbf{diag}(\widehat{\mathcal{A}}_{\mathbf{j},\mathbf{k}}).$ (4)
By Definition 2.2 and the identities (2) and (3), we have
$\textbf{Vec}(\mathcal{C})$ $=\textbf{Vec}(\mathcal{A}\ast_{Ht}\mathcal{B})$
$=(\textbf{circ}(\mathcal{A}_{1,\mathbf{i}})+\mathbf{j}\textbf{circ}(\mathcal{A}_{\mathbf{j},\mathbf{k}}))\textbf{Vec}(\mathcal{B})$
$=(F_{n_{3}}^{\ast}\otimes I_{n_{1}})[\textbf{diag}(\widehat{\mathcal{A}}_{1,\mathbf{i}})+\mathbf{j}\textbf{diag}(\widehat{\mathcal{A}}_{\mathbf{j},\mathbf{k}})](F_{n_{3}}\otimes I_{n_{2}})\textbf{Vec}(\mathcal{B})$
$=\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})[\textbf{diag}(\widehat{\mathcal{A}}_{1,\mathbf{i}})+\mathbf{j}\textbf{diag}(\widehat{\mathcal{A}}_{\mathbf{j},\mathbf{k}})]\textbf{Vec}(\widehat{\mathcal{B}}).$
Thus
$\textbf{Vec}(\widehat{\mathcal{C}})=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{C})=[\textbf{diag}(\widehat{\mathcal{A}}_{1,\mathbf{i}})+\mathbf{j}\textbf{diag}(\widehat{\mathcal{A}}_{\mathbf{j},\mathbf{k}})]\textbf{Vec}(\widehat{\mathcal{B}}).$
Combing with equation (4) yields $\textbf{Vec}(\widehat{\mathcal{C}})=\textbf{diag}(\widehat{\mathcal{A}})\textbf{Vec}(\widehat{\mathcal{B}})$. Therefore,
$\textbf{diag}(\widehat{\mathcal{C}})=\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}}).$
$\blacksquare$
According to the above result, we can give an equivalent definition of the Ht-product of two third-order reduced biquaternion tensors.
Definition 2.4
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$. The $Ht$-product of $\mathcal{A}$ and $\mathcal{B}$ is defined as
$\mathcal{A}\ast_{Ht}\mathcal{B}={\textbf{Fold}}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}})(e\otimes I_{n_{4}})\right),$
where $e$ is an $n_{3}\times 1$ vector with all the entries equal to 1.
This Ht-product obeys some usual algebraic operation laws, such as the associative law and the distributive law.
Proposition 2.5
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}},\ \mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$ and $\mathcal{C}\in\mathbb{H}_{c}^{n_{4}\times n_{5}\times n_{3}}$. Then
$(\mathcal{A}\ast_{Ht}\mathcal{B})\ast_{Ht}\mathcal{C}=\mathcal{A}\ast_{Ht}(\mathcal{B}\ast_{Ht}\mathcal{C}).$
Proof.  By Definition 2.4, we have
$(\mathcal{A}\ast_{Ht}\mathcal{B})\ast_{Ht}\mathcal{C}=$ $\textbf{Fold}(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}})(e\otimes I_{n_{4}}))\ast_{Ht}\mathcal{C}$
$=$ $\textbf{Fold}(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})(\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}}))\textbf{diag}(\widehat{\mathcal{C}})(e\otimes I_{n_{5}})),$
and
$\mathcal{A}\ast_{Ht}(\mathcal{B}\ast_{Ht}\mathcal{C})=$ $\mathcal{A}\ast_{Ht}\textbf{Fold}(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{2}})\textbf{diag}(\widehat{\mathcal{B}})\textbf{diag}(\widehat{\mathcal{C}})(e\otimes I_{n_{5}}))$
$=$ $\textbf{Fold}(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})\textbf{diag}(\widehat{\mathcal{A}})(\textbf{diag}(\widehat{\mathcal{B}})\textbf{diag}(\widehat{\mathcal{C}}))(e\otimes I_{n_{5}})).$
Notice that $\textbf{diag}(\widehat{\mathcal{A}}),\textbf{diag}(\widehat{\mathcal{B}})$ and $\textbf{diag}(\widehat{\mathcal{C}})$ are matrices over $\mathbb{H}_{c}$. Hence the associative law holds, and thus the result holds. $\blacksquare$
In a similar way as in Proposition 2.5, we have the following distributive laws. For brevity, we omit the proofs.
Proposition 2.6
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}},\ \mathcal{B},\ \mathcal{C}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$ and $\mathcal{D}\in\mathbb{H}_{c}^{n_{4}\times n_{5}\times n_{3}}$. Then
$\mathcal{A}\ast_{Ht}(\mathcal{B}+\mathcal{C})=\mathcal{A}\ast_{Ht}\mathcal{B}+\mathcal{A}\ast_{Ht}\mathcal{C},$
$(\mathcal{B}+\mathcal{C})\ast_{Ht}\mathcal{D}=\mathcal{B}\ast_{Ht}\mathcal{D}+\mathcal{C}\ast_{Ht}\mathcal{D}.$
Suppose $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$. Using a similar approach in [13], we calculate the conjugate transpose $\mathcal{A}^{\ast}\in\mathbb{C}^{n_{2}\times n_{1}\times n_{3}}$ of $\mathcal{A}$ as follows: conjugately transposing each frontal slice of $\mathcal{A}$ first, and reversing the order of conjugately transposed frontal slices $2$ through $n_{3}$ next.
Definition 2.7
The conjugate transpose $\mathcal{A}^{\ast}$ of $\mathcal{A}=\mathcal{A}_{1,\mathbf{i}}+\mathbf{j}\mathcal{A}_{j,k}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ with $\mathcal{A}_{1,\mathbf{i}},\mathcal{A}_{\mathbf{j},\mathbf{k}}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ is defined as $\mathcal{A}^{\ast}=\mathcal{A}_{1,\mathbf{i}}^{\ast}+j\mathcal{A}_{\mathbf{j},\mathbf{k}}^{\ast}$ with $\mathcal{A}_{1,\mathbf{i}}^{\ast},\mathcal{A}_{\mathbf{j},\mathbf{k}}^{\ast}\in\mathbb{C}^{n_{2}\times n_{1}\times n_{3}}$.
It is not difficult to verify that $(\mathcal{A}^{\ast})^{\ast}=\mathcal{A}$ and $(\mathcal{A}+\mathcal{B})^{\ast}=\mathcal{A}^{\ast}+\mathcal{B}^{\ast}$ for $\mathcal{A},\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$.
Definition 2.8
The identity tensor $\mathcal{I}_{nnm}\in\mathbb{H}_{c}^{n\times n\times m}$ (simply denoted by $\mathcal{I}$) is the tensor in which the first frontal slice is $I_{n}$ and the other slice matrices are zero.
We can obtain that $\mathcal{I}\ast_{Ht}\mathcal{I}=\mathcal{I}$$\mathcal{I}^{\ast}=\mathcal{I}$ and $\mathcal{A}\ast_{Ht}\mathcal{I}=\mathcal{I}\ast_{Ht}\mathcal{A}=\mathcal{A}.$
Definition 2.9
A tensor $\mathcal{A}\in\mathbb{H}_{c}^{n\times n\times m}$ is called invertible if there is a tensor $\mathcal{B}\in\mathbb{H}_{c}^{n\times n\times m}$ such that $\mathcal{A}\ast_{Ht}\mathcal{B}=\mathcal{I}$ and $\mathcal{B}\ast_{Ht}\mathcal{A}=\mathcal{I}$.
By using Proposition 2.5 and the fact $\mathcal{A}\ast_{Ht}\mathcal{I}=\mathcal{I}\ast_{Ht}\mathcal{A}=\mathcal{A}$, we have the following result.
Proposition 2.10
The inverse of a reduced biquaternion tensor $\mathcal{A}\in\mathbb{H}_{c}^{n\times n\times m}$ is unique if it exists, denoted by $\mathcal{A}^{-1}$.
Definition 2.11
Let $\mathcal{A}\in\mathbb{H}_{c}^{n\times n\times m}$. $\mathcal{A}$ is called idempotent if $\mathcal{A}\ast_{Ht}\mathcal{A}=\mathcal{A}$; $\mathcal{A}$ is called unitary (resp. Hermitian) if $\mathcal{A}\ast_{Ht}{\mathcal{A}}^{\ast}={\mathcal{A}}^{\ast}\ast_{Ht}\mathcal{A}=\mathcal{I}\ (resp.\ \mathcal{A}={\mathcal{A}}^{\ast}).$
Lemma 2.12
Let $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$. Then $\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$
Proof.  It is easy to see that $\textbf{circ}({\mathcal{A}}^{\ast})=(\textbf{circ}({\mathcal{A}}))^{\ast}$ for $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}.$ According to the identity (2), we have
$\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})$ $=(F_{n_{3}}\otimes I_{n_{2}})\textbf{circ}({\mathcal{A}}^{\ast})(F_{n_{3}}^{\ast}\otimes I_{n_{1}})$
$=(F_{n_{3}}\otimes I_{n_{2}})(\textbf{circ}({\mathcal{A}}))^{\ast}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})$
$=(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$
$\blacksquare$
The above result over $\mathbb{C}$ can be extended to $\mathbb{H}_{c}$ as follows.
Proposition 2.13
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then $\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$
Proof.  Since $\widehat{{\mathcal{A}}^{\ast}}$ is the DFT of ${\mathcal{A}}^{\ast}$ and $F_{n_{3}}\mathbf{j}=\mathbf{j}F_{n_{3}}$, it follows from (3) that
$\textbf{Vec}(\widehat{{\mathcal{A}}^{\ast}})$ $=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{2}})\textbf{Vec}({\mathcal{A}}^{\ast})$
$=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{2}})(\textbf{Vec}({\mathcal{A}}_{1,\mathbf{i}}^{\ast})+\mathbf{j}\textbf{Vec}({\mathcal{A}}_{\mathbf{j},\mathbf{k}}^{\ast}))$
$=\textbf{Vec}(\widehat{{\mathcal{A}}_{1,\mathbf{i}}^{\ast}})+\mathbf{j}\textbf{Vec}(\widehat{{\mathcal{A}}_{\mathbf{j},\mathbf{k}}^{\ast}}),$
which is equivalent to
$\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=\textbf{diag}(\widehat{{\mathcal{A}}_{1,\mathbf{i}}^{\ast}})+\mathbf{j}\textbf{diag}(\widehat{{\mathcal{A}}_{\mathbf{j},\mathbf{k}}^{\ast}}).$ (5)
Similarly, we have $\textbf{diag}(\widehat{\mathcal{A}})=\textbf{diag}(\widehat{{\mathcal{A}}_{1,\mathbf{i}}})+\mathbf{j}\textbf{diag}(\widehat{{\mathcal{A}}_{\mathbf{j},\mathbf{k}}})\in\mathbb{H}_{c}^{n_{3}n_{1}\times n_{3}n_{2}}.$ Its conjugate transpose is $(\textbf{diag}(\hat{\mathcal{A}}))^{\ast}=(\textbf{diag}(\widehat{{\mathcal{A}}_{1,\mathbf{i}}}))^{\ast}+\mathbf{j}(\textbf{diag}(\widehat{{\mathcal{A}}_{\mathbf{j},\mathbf{k}}}))^{\ast}.$ Combing with the identity (5) and Lemma 2.12, we have $\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$blacksquare
Corollary 2.14
$\mathcal{A}\in\mathbb{H}_{c}^{n\times n\times m}$ is Hermitian (resp. unitary) if and only if $\textbf{diag}(\widehat{\mathcal{A}})$ is Hermitian (resp. unitary).
Proof.  By Definition 2.8, it is clear that $\textbf{diag}(\widehat{\mathcal{I}})=I_{nm}.$ Then, by Theorem 2.3 and Proposition 2.13, the result follows. $\blacksquare$
Proposition 2.15
Given $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$, we have $(\mathcal{A}\ast_{Ht}\mathcal{B})^{\ast}=\mathcal{B}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}.$
Proof.  From Theorem 2.3 and Proposition 2.13, we only need to show
$(\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}}))^{\ast}=(\textbf{diag}(\widehat{\mathcal{B}}))^{\ast}(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$
In fact, this equality holds for matrices $\textbf{diag}(\widehat{\mathcal{A}})$ and $\textbf{diag}(\widehat{\mathcal{B}})$ over $\mathbb{H}_{c}$. $\blacksquare$
3 The Ht-SVD of third-order tensors over $\mathbb{H}_{c}$
The SVD of a matrix over $\mathbb{H}_{c}$ has been discussed in some papers (see, e.g., [11]). Let $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times n}$. If $A_{(c),1}$ and $A_{(c),2}$ have the following SVDs:
$A_{(c),1}=U_{1}\Sigma_{1}V_{1}^{\ast}$ and $A_{(c),2}=U_{2}\Sigma_{2}V_{2}^{\ast},$
then the SVD of $A$ is
$A=U\Sigma V^{\ast},$ (6)
where $U=U_{1}e_{1}+U_{2}e_{2}$ and $V=V_{1}e_{1}+V_{2}e_{2}$ are unitary matrices, $\Sigma=\Sigma_{1}e_{1}+\Sigma_{2}e_{2}$ is diagonal but not a real matrix in general. We define the rank of a matrix $A\in\mathbb{H}_{c}^{m\times n}$ is the number of its non-zero singular values.
Recall that a third-order tensor is called $f$-diagonal if each frontal slice is diagonal ([5]). Next, we propose the Ht-SVD of a reduced biquaternion tensor.
Theorem 3.1
(Ht-SVD) Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then there exist unitary tensors $\mathcal{U}\in\mathbb{H}_{c}^{n_{1}\times n_{1}\times n_{3}}$ and $\mathcal{V}\in\mathbb{H}_{c}^{n_{2}\times n_{2}\times n_{3}}$ such that
$\mathcal{A}=\mathcal{U}\ast_{Ht}\mathcal{S}\ast_{Ht}\mathcal{V}^{\ast},$
where $\mathcal{S}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ is $f$-diagonal. Such decomposition is called $Ht$-SVD of $\mathcal{A}$.
Proof.  We first transform $\mathcal{A}$ into $\widehat{\mathcal{A}}$ by using DFT along the third mode, and then obtain the block diagonal matrix $\textbf{diag}(\widehat{\mathcal{A}})$ by (1). In addition, we consider the SVD of each $\widehat{A}^{(s)}$ in $\textbf{diag}(\widehat{\mathcal{A}})$ with formula (6). Suppose $\widehat{A}^{(s)}=\widehat{U}^{(s)}\widehat{\Sigma}^{(s)}(\widehat{V}^{(s)})^{\ast},s=1,2,\cdots,n_{3}.$ Then we have
$\textbf{diag}(\widehat{\mathcal{A}})=\begin{bmatrix}\widehat{U}^{(1)}&&&\\ &\widehat{U}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{U}^{(n_{3})}\\ \end{bmatrix}\begin{bmatrix}\widehat{\Sigma}^{(1)}&&&\\ &\widehat{\Sigma}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{\Sigma}^{(n_{3})}\\ \end{bmatrix}\begin{bmatrix}(\widehat{V}^{(1)})^{\ast}&&&\\ &(\widehat{V}^{(2)})^{\ast}&&\\ &&\ddots&\\ &&&(\widehat{V}^{(n_{3})})^{\ast}\\ \end{bmatrix}.$
Setting
$\textbf{diag}(\widehat{\mathcal{U}})=\begin{bmatrix}\widehat{U}^{(1)}&&&\\ &\widehat{U}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{U}^{(n_{3})}\\ \end{bmatrix},\quad\textbf{diag}(\widehat{\mathcal{V}})=\begin{bmatrix}\widehat{V}^{(1)}&&&\\ &\widehat{V}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{V}^{(n_{3})}\\ \end{bmatrix},$
and
$\textbf{diag}(\hat{\mathcal{S}})=\begin{bmatrix}\widehat{\Sigma}^{(1)}&&&\\ &\widehat{\Sigma}^{(2)}&&\\ &&\ddots&\\ &&&\widehat{\Sigma}^{(n_{3})}\\ \end{bmatrix},$
we have
$\textbf{diag}(\widehat{\mathcal{A}})=\textbf{diag}(\widehat{\mathcal{U}})\textbf{diag}(\widehat{\mathcal{S}})(\textbf{diag}(\widehat{\mathcal{V}}))^{\ast},$ (7)
in which $\textbf{diag}(\widehat{\mathcal{U}})$ and $\textbf{diag}(\widehat{\mathcal{V}})$ are unitary matrices, and $\textbf{diag}(\widehat{\mathcal{S}})$ is a block diagonal matrix with elements being diagonal matrices in general. Next, we construct reduced biquaternion tensors $\mathcal{U},\mathcal{V}$ and $\mathcal{S}$ as follows:
$\mathcal{U}\doteq\textbf{Fold}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})\textbf{diag}(\widehat{\mathcal{U}})(e\otimes I_{n_{1}})\right),$
$\mathcal{V}\doteq\textbf{Fold}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{2}})\textbf{diag}(\widehat{\mathcal{V}})(e\otimes I_{n_{2}})\right),$
$\mathcal{S}\doteq\textbf{Fold}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})\textbf{diag}(\widehat{\mathcal{S}})(e\otimes I_{n_{2}})\right).$
Observing that $\textbf{Vec}(\widehat{\mathcal{U}})=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{U}),\textbf{Vec}(\widehat{\mathcal{V}})=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{2}})\textbf{Vec}(\mathcal{V}),\textbf{Vec}(\widehat{\mathcal{S}})=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{S}).$ That is, $\widehat{\mathcal{U}},\widehat{\mathcal{V}},\widehat{\mathcal{S}}$ are the DFT of tensors $\mathcal{U},\mathcal{V},\mathcal{S}$, respectively. Thus by applying Theorem 2.3 and Proposition 2.13 , we infer from the equation (7) that
$\mathcal{A}=\mathcal{U}\ast_{Ht}\mathcal{S}\ast_{Ht}\mathcal{V}^{\ast}.$
Obviously, $\mathcal{S}$ is $f$-diagonal. And according to Corollary 2.14, we know that $\mathcal{U}$, $\mathcal{V}$ are our required unitary tensors. $\blacksquare$
Based on Theorem 3.1, we give the reduced biquaternion tensor SVD decomposition (Ht-SVD) algorithm as follows. In MATLAB command, for the complex tensor $\mathcal{C}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$, the DFT of $\mathcal{C}$ and its inverse function are denoted as:
$\hat{\mathcal{C}}=\text{fft}(\mathcal{C},\left[\,\right],3),\ \ \mathcal{C}=\text{ifft}(\hat{\mathcal{C}},\left[\,\right],3).$
Algorithm 1 (SVD-RBT) SVD for a Reduced Biquaternion Tensor
1:  Input: $\mathcal{A}_{(c),1},\mathcal{A}_{(c),2}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ from given $\mathcal{A}=\mathcal{A}_{(c),1}e_{1}+\mathcal{A}_{(c),2}e_{2}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$
2:  Output: $\mathcal{U}\in\mathbb{H}_{c}^{n_{1}\times n_{1}\times n_{3}},\mathcal{S}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{V}\in\mathbb{H}_{c}^{n_{2}\times n_{2}\times n_{3}},$ such that
$\mathcal{A}=\mathcal{U}\ast_{Ht}\mathcal{S}\ast_{Ht}\mathcal{V}^{\ast}$
3:  $\hat{\mathcal{A}}_{(c),1}$=fft$(\mathcal{A}_{(c),1},[],3)$;$\hat{\mathcal{A}}_{(c),2}$=fft$(\mathcal{A}_{(c),2},[],3)$
4:  Calculate each frontal slice of $\hat{\mathcal{U}},\hat{\mathcal{S}}$ and $\hat{\mathcal{V}}$ from $\hat{\mathcal{A}}$:for $i=1,...n_{3},$ do$[\hat{U}_{1}^{(i)},\hat{S}_{1}^{(i)},\hat{V}_{1}^{(i)}]=svd(\hat{\mathcal{A}}_{(c),1}^{(i)})$, $[\hat{U}_{2}^{(i)},\hat{S}_{2}^{(i)},\hat{V}_{2}^{(i)}]=svd(\hat{\mathcal{A}}_{(c),2}^{(i)})$end for
[…]
6:  $\mathcal{U}=\mathcal{U}_{1}e_{1}+\mathcal{U}_{2}e_{2};$ $\mathcal{S}=\mathcal{S}_{1}e_{1}+\mathcal{S}_{2}e_{2};$ $\mathcal{V}=\mathcal{V}_{1}e_{1}+\mathcal{V}_{2}e_{2};$
7:  return $\mathcal{U}$, $\mathcal{S}$, $\mathcal{V}$
Complexity Analysis of Algorithm 1: For a given reduced biquaternion tensor $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$, the computational complexity of each iteration of Algorithm 1 (SVD-RBT) is primarily driven by two major components: the fast fourier transform (FFT) and its inverse, and the SVD of the tensor’s frontal slices. The computational complexity of FFT and inverse FFT operations is $O(n_{1}n_{2}n_{3}\log n_{3})$. The SVD process, applied separately to each frontal slice, has a complexity of $O(n_{3}\min(n_{1}^{2}n_{2},n_{1}n_{2}^{2}))$. Additionally, the tensor reconstruction step, which involves simple element-wise addition, has a complexity of $O(n_{1}n_{2}n_{3})$. Thus, the total computational complexity for each iteration of the algorithm is $O(n_{1}n_{2}n_{3}\log n_{3}+n_{3}\min(n_{1}^{2}n_{2},n_{1}n_{2}^{2}))$.
4 The MP inverses of reduced biquaternion tensors
We refer the readers to [1] for Moore-Penrose (MP) inverses. In this section, as an application of Ht-SVD of a RB tensor, we define the MP inverse for a RB tensor and find a formula for the MP inverse by using Theorem 3.1. Moreover, we explore some theoretical properties of the MP inverses of RB tensors.
For a given matrix $A$, its MP inverse is the unique matrix $X$ (denoted as $A^{\dagger}$) satisfying
$AXA=A,\ XAX=X,\ (AX)^{\ast}=AX,\ (XA)^{\ast}=XA.$
According to the properties of reduced biquaternion matrices, we can easily derive the MP inverse of a reduced biquaternion matrix as follows.
Theorem 4.1
Let $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times n},$ where $A_{(c),1},\ A_{(c),2}\in\mathbb{C}^{m\times n}$. The MP inverse of $A$ is $A^{\dagger}=A_{(c),1}^{\dagger}e_{1}+A_{(c),2}^{\dagger}e_{2}\in\mathbb{H}_{c}^{n\times m}.$
Now, we introduce the definition of the MP inverse of a third-order reduced biquaternion tensor.
Definition 4.2
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. If there exists $\mathcal{X}\in\mathbb{H}_{c}^{n_{2}\times n_{1}\times n_{3}}$ satisfying:
  • (i)
    $\mathcal{A}\ast_{Ht}\mathcal{X}\ast_{Ht}\mathcal{A}=\mathcal{A};$
  • (ii)
    $\mathcal{X}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{X};$
  • (iii)
    $(\mathcal{A}\ast_{Ht}\mathcal{X})^{\ast}=\mathcal{A}\ast_{Ht}\mathcal{X};$
  • (iv)
    $(\mathcal{X}\ast_{Ht}\mathcal{A})^{\ast}=\mathcal{X}\ast_{Ht}\mathcal{A},$
then $\mathcal{X}$ is said to be an MP inverse of $\mathcal{A}$, denoted by $\mathcal{A}^{\dagger}.$
For each $n_{1}\times n_{2}\times n_{3}$ reduced biquaternion tensor, the following theorem shows that it has a unique MP inverse.
Theorem 4.3
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and its Ht-SVD be $\mathcal{A}=\mathcal{U}\ast_{Ht}\mathcal{S}\ast_{Ht}\mathcal{V}^{\ast},$ where $\mathcal{U}\in\mathbb{H}_{c}^{n_{1}\times n_{1}\times n_{3}},\mathcal{V}\in\mathbb{H}_{c}^{n_{2}\times n_{2}\times n_{3}}$ are unitary and $\mathcal{S}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ is f-diagonal. Then $\mathcal{A}$ has a unique MP inverse
$\mathcal{A}^{\dagger}=\mathcal{V}\ast_{Ht}\mathcal{S}^{\dagger}\ast_{Ht}\mathcal{U}^{\ast},$ (8)
where $\mathcal{S}^{\dagger}=\textbf{Fold}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{2}})(\textbf{diag}(\widehat{\mathcal{S}}))^{\dagger}(e\otimes I_{n_{1}})\right)$.
Proof.  From Theorem 2.3, Proposition 2.5, Proposition 2.15 and Definition 4.1, it is not difficult to verify that $\textbf{diag}(\widehat{\mathcal{S}^{\dagger}})=(\textbf{diag}(\widehat{\mathcal{S}}))^{\dagger}$ and the expression of $\mathcal{A}^{\dagger}$ given in (8) satisfies all four conditions in Definition 4.2. Next, we need to prove the uniqueness of the MP inverse. Suppose that $\mathcal{X}$ and $\mathcal{Y}$ are two MP inverses of $\mathcal{A}$, respectively. Then
$\mathcal{X}$ $=\mathcal{X}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{X}\ast_{Ht}\mathcal{X}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}$
$=\mathcal{X}\ast_{Ht}\mathcal{X}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}\mathcal{Y}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}$
$=\mathcal{X}\ast_{Ht}\mathcal{Y}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{X}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Y}$
$=\mathcal{X}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Y}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Y}$
$=(\mathcal{X}\ast_{Ht}\mathcal{A})^{\ast}\ast_{Ht}(\mathcal{Y}\ast_{Ht}\mathcal{A})^{\ast}\ast_{Ht}\mathcal{Y}$
$=\mathcal{A}^{\ast}\ast_{Ht}\mathcal{Y}^{\ast}\ast_{Ht}\mathcal{Y}=\mathcal{Y}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Y}$
$=\mathcal{Y}.$
The proof is completed. $\blacksquare$
Next, we explore some properties of the MP inverse of a reduced biquaternion tensor $\mathcal{A}$.
Proposition 4.4
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then
  • (a)
    $(\mathcal{A}^{\dagger})^{\dagger}=\mathcal{A}$;
  • (b)
    $(\mathcal{A}^{\dagger})^{\ast}=(\mathcal{A}^{\ast})^{\dagger}$;
  • (c)
    $(\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast})^{\dagger}=(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}\mathcal{A}^{\dagger}$ and $(\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A})^{\dagger}=\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}^{\ast})^{\dagger}$;
  • (d)
    $(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})^{\dagger}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}$ ( $\mathcal{P}$ and $\mathcal{Q}$ are unitary reduced biquaternion tensors );
  • (e)
    If $\mathcal{A}=(a_{i_{1}i_{2}i_{3}})\in\mathbb{H}_{c}^{n_{1}\times n_{1}\times n_{3}}$ is f-diagonal, then $\mathcal{A}^{\dagger}$ is f-diagonal, and $(\mathcal{A}^{\dagger}_{i_{1}i_{2}i_{3}})=(a_{i_{1}i_{2}i_{3}})^{\dagger}$. In this case, if each diagonal entry in each frontal slice has an inverse, then $\mathcal{A}^{\dagger}=\mathcal{A}^{-1}$;
  • (f)
    $\mathcal{A}^{\dagger}=\mathcal{A}^{\ast}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast})^{\dagger}=(\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A})^{\dagger}\ast_{Ht}\mathcal{A}^{\ast}$;
  • (g)
    $\mathcal{A}=\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}(\mathcal{A}^{\ast})^{\dagger}=(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}$;
  • (h)
    $\mathcal{A}^{\ast}=\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}$;
  • (i)
    $\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}=(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}\mathcal{A}^{\ast}$ and $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}=\mathcal{A}^{\ast}\ast_{Ht}(\mathcal{A}^{\ast})^{\dagger}$;
  • (j)
    $\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}=(\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast})^{\dagger}$;
  • (k)
    $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}=(\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A})^{\dagger}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}=\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}\ast_{Ht}(\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A})^{\dagger}$;
  • (l)
    $\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}$ if $\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}$;
  • (m)
    $\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}$ and $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}$ are idempotent;
  • (n)
    If $\mathcal{A}$ is Hermitian and idempotent, then $\mathcal{A}^{\dagger}=\mathcal{A}$.
Proof.  We only prove (b), (c), (d), (l) and (n). In the analogous way, one can obtain the remaining properties.
  • (b)
    Using Propositions 2.5 and 2.15, one can obtain that $(\mathcal{A}^{\dagger})^{\ast}$ is the MP inverse of $\mathcal{A}^{\ast}$ by direct calculations.
  • (c)
    By (b), Proposition 2.5, Proposition 2.15 and the fact $(\mathcal{A}^{\ast})^{\ast}=\mathcal{A}$, one can get that $(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}\mathcal{A}^{\dagger}$ is the MP inverse of $\mathcal{A}\ast_{Ht}\mathcal{A}^{\ast}$ and $\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}^{\ast})^{\dagger}$ is the MP inverse of $\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}$.
  • (d)
    Suppose $\mathcal{P}\ast_{Ht}\mathcal{P}^{\ast}=\mathcal{P}^{\ast}\ast_{Ht}\mathcal{P}=\mathcal{I}$ and $\mathcal{Q}\ast_{Ht}\mathcal{Q}^{\ast}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{Q}=\mathcal{I}$. Then we have
    $\begin{array}[]{lll}&(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})\ast_{Ht}(\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast})\ast_{Ht}(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})\\ &=\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q}=\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q},\end{array}$
    $\begin{array}[]{lll}&(\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast})\ast_{Ht}(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})\ast_{Ht}(\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast})\\ &=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast},\end{array}$
    $\begin{array}[]{lll}&(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q}\ast_{Ht}\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast})^{\ast}\\ &=\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}=\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q}\ast_{Ht}\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast},\end{array}$
    and
    $(\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}\ast_{Ht}\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})^{\ast}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}\ast_{Ht}\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q}.$
    Hence, by Definition 4.2 we can derive that $(\mathcal{P}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{Q})^{\dagger}=\mathcal{Q}^{\ast}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{P}^{\ast}.$
  • (l)
    It follows from (i) and (j).
  • (n)
    If $\mathcal{A}$ is Hermitian and idempotent, then $\mathcal{A}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{A}=\mathcal{A}$ and $(\mathcal{A}\ast_{Ht}\mathcal{A})^{\ast}=\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{A}\ast_{Ht}\mathcal{A}$. Thus by Definition 4.2 we have $\mathcal{A}^{\dagger}=\mathcal{A}$.
$\blacksquare$
In general, $(\mathcal{A}\ast_{Ht}\mathcal{B})^{\dagger}\neq\mathcal{B}^{\dagger}\ast_{Ht}\mathcal{A}^{\dagger}$. Under some conditions, we have the equality case given as follows.
Proposition 4.5
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$. Then $(\mathcal{A}\ast_{Ht}\mathcal{B})^{\dagger}=\mathcal{B}^{\dagger}\ast_{Ht}\mathcal{A}^{\dagger}$ if one of the following conditions satisfies.
(a) $\mathcal{A}=\mathcal{B}^{\dagger}$;   (b) $\mathcal{A}=\mathcal{B}^{\ast}$;  (c) $\mathcal{A}^{\ast}\ast_{Ht}\mathcal{A}=\mathcal{I}$;  (d) $\mathcal{B}\ast_{Ht}\mathcal{B}^{\ast}=\mathcal{I}$.
Denote $\mathcal{L}_{\mathcal{A}}=\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}\mathcal{A},\ \mathcal{R}_{\mathcal{A}}=\mathcal{I}-\mathcal{A}\ast_{Ht}{\mathcal{A}}^{\dagger}.$ $\mathcal{O}$ represents zero tensor with all the entries being zero. Obviously, $\mathcal{A}\ast_{Ht}\mathcal{O}=\mathcal{O}\ast_{Ht}\mathcal{A}=\mathcal{O}.$
Proposition 4.6
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then
  • (a)
    $\mathcal{A}\ast_{Ht}\mathcal{L}_{\mathcal{A}}=\mathcal{O}$ and $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{A}=\mathcal{O}$;
  • (b)
    $\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{A}^{\dagger}=\mathcal{O}$ and $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{R}_{\mathcal{A}}=\mathcal{O}$;
  • (c)
    $\mathcal{L}_{\mathcal{A}^{\ast}}=(\mathcal{R}_{\mathcal{A}})^{\ast}=\mathcal{R}_{\mathcal{A}}$ and $\mathcal{R}_{\mathcal{A}^{\ast}}=(\mathcal{L}_{\mathcal{A}})^{\ast}=\mathcal{L}_{\mathcal{A}}$;
  • (d)
    $\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{L}_{\mathcal{A}}=\mathcal{L}_{\mathcal{A}}$ and $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{R}_{\mathcal{A}}=\mathcal{R}_{\mathcal{A}}$;
  • (e)
    $(\mathcal{L}_{\mathcal{A}})^{\dagger}=\mathcal{L}_{\mathcal{A}}$ and $(\mathcal{R}_{\mathcal{A}})^{\dagger}=\mathcal{R}_{\mathcal{A}}$;
  • (f)
    $\mathcal{L}_{\mathcal{A}}=\mathcal{L}_{\mathcal{A}^{\ast}\ast_{Ht}{\mathcal{A}}}$ and $\mathcal{R}_{\mathcal{A}}=\mathcal{R}_{\mathcal{A}\ast_{Ht}{\mathcal{A}^{\ast}}}$.
Proof.  By straightforward calculations, one can prove that (a), (b) and (d) hold.
For (c), the first equality follows from
$\mathcal{L}_{\mathcal{A}^{\ast}}=\mathcal{I}-(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}{\mathcal{A}^{\ast}}=\mathcal{I}-(\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger})^{\ast}=(\mathcal{I}-\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger})^{\ast}=(\mathcal{R}_{\mathcal{A}})^{\ast}$
and
$\mathcal{L}_{\mathcal{A}^{\ast}}=\mathcal{I}-(\mathcal{A}^{\ast})^{\dagger}\ast_{Ht}{\mathcal{A}^{\ast}}=\mathcal{I}-(\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger})^{\ast}=\mathcal{I}-\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger}=\mathcal{R}_{\mathcal{A}}.$
In an analogous manner, one can prove the second part.
For (e), it follows immediately from (c), (d) and Proposition 4.4.
For (f), $\mathcal{L}_{\mathcal{A}^{\ast}\ast_{Ht}{\mathcal{A}}}=\mathcal{I}-(\mathcal{A}^{\ast}\ast_{Ht}{\mathcal{A}})^{\dagger}\ast_{Ht}(\mathcal{A}^{\ast}\ast_{Ht}{\mathcal{A}})=\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}({\mathcal{A}}^{\dagger})^{\ast}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}{\mathcal{A}}=\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}{\mathcal{A}}\ast_{Ht}\mathcal{A}^{\dagger}\ast_{Ht}{\mathcal{A}}=\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}{\mathcal{A}}=\mathcal{L}_{\mathcal{A}}.$ The second part can be proved in a similar way. $\blacksquare$
5 Reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$
We mainly discuss the consistency of RB tensor equation
$\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$ (9)
by using the MP inverse demonstrated in Section 4. The solvability conditions are found and a general expression of the solutions is provided when the solutions exist. Except that, we derive the least-squares solutions/the minimal norm least-squares solution to (9). Furthermore, we also derive solvability conditions and the general solutions involving Hermicity. One of the applications of solving this tensor equations is color video deblurring given in Section 6.
Theorem 5.1
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}.$ Then equation (9) has solutions if and only if $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}$. Furthermore, the general solutions to equation (9) can be written as
$\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{Y},$ (10)
with $\mathcal{Y}$ be any reduced biquaternion tensor with compatible size.
Proof.  Assume that $\mathcal{C}$ is a solution to reduced biquaternion tensor equation (9). Then by Proposition 2.5 and Proposition 4.6, we have
$\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{C}=\mathcal{O}.$
Conversely, if $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}$, it is easy to see that $\mathcal{X}_{0}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}$ is a solution to equation (9), and so is $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{Y}$.
Next we show each solution $\mathcal{X}$ of equation (9) has the form of $\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{X}$. Note that $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$, thus $\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}-\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{X}+\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}\mathcal{A})\ast_{Ht}\mathcal{X},$ as required. Based on the above, we can conclude that (10) is the general expression of the solutions. $\blacksquare$
Theorem 5.2
Let $\mathcal{A},\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then equation (9) has a Hermitian solution if and only if $\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast}=\mathcal{B}\ast_{Ht}\mathcal{A}^{\ast}$ and $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}.$ Moreover, if (9) is solvable, then the general Hermitian solutions to the reduced biquaternion tensor equation (9) can be written as
$\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B})^{\ast}-\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast})\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{U}\ast_{Ht}\mathcal{L}_{\mathcal{A}},$ (11)
where $\mathcal{U}=\mathcal{U}^{\ast}$ is an arbitrary tensor with appropriate size.
Proof.  If equation (9) has a Hermitian solution $\mathcal{C}$, then by Propositions 2.5, 2.15 and Definition 4.2, we obtain
$\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast}=\mathcal{A}\ast_{Ht}\mathcal{C}^{\ast}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{A}\ast_{Ht}\mathcal{C}\ast_{Ht}\mathcal{A}^{\ast}=\mathcal{B}\ast_{Ht}\mathcal{A}^{\ast},$
and $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{C}=\mathcal{O}.$
On the other hand, assume that $\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast}=\mathcal{B}\ast_{Ht}\mathcal{A}^{\ast}$ and $R_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}$ hold. Then by applying Propositions 2.5 and 2.15, we get that
$\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B})^{\ast}-\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast})\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{U}\ast_{Ht}\mathcal{L}_{\mathcal{A}}$
is a Hermitian solution to reduced biquaternion tensor equation (9).
Next we show that each solution of reduced biquaternion tensor equation (9) should be in the form of (11). Assume $\mathcal{X}_{0}$ is an arbitrary Hermitian solution to tensor equation (9). Setting $\mathcal{U}=\mathcal{X}_{0}$ yields
$\mathcal{X}=$ $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B})^{\ast}-\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast})\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{X}_{0}\ast_{Ht}\mathcal{L}_{\mathcal{A}}$
$=$ $\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B})^{\ast}-\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast})\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}$
$+(\mathcal{X}_{0}-\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{X}_{0})\ast_{Ht}(\mathcal{I}-\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A})$
$=$ $(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B})^{\ast}-\mathcal{A}^{\dagger}\ast_{Ht}(\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast})\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}$
$+\mathcal{X}_{0}-\mathcal{X}_{0}\ast_{Ht}(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A})+\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}\ast_{Ht}\mathcal{A}^{\ast}\ast_{Ht}(\mathcal{A}^{\dagger})^{\ast}$
$=$ $(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{X}_{0})^{\ast}+\mathcal{X}_{0}-\mathcal{X}_{0}\ast_{Ht}(\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A})^{\ast}$
$=$ $\mathcal{X}_{0},$
which implies that each Hermitian solution to reduced biquaternion tensor equation (9) can be represented as (11). Thus, (11) is the general expression of Hermitian solutions to equation (9). $\blacksquare$
Next, we deal with the least-squares problem of RB tensor equation (9). For a tensor $\mathcal{A}=\mathcal{A}_{1,\mathbf{i}}+\mathbf{j}\mathcal{A}_{\mathbf{j},\mathbf{k}}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ with $\mathcal{A}_{1,\mathbf{i}},\mathcal{A}_{\mathbf{j},\mathbf{k}}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$, we define its Frobenius norm as
$\|\mathcal{A}\|_{F}^{2}=\|\textbf{Vec}(\mathcal{A})\|_{F}^{2}=\|\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})\|_{F}^{2}+\|\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}})\|_{F}^{2}.$ (12)
where $\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})=(c_{ij})_{n_{1}n_{3}\times{n_{2}}},$ $\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}})=(d_{ij})_{n_{1}n_{3}\times{n_{2}}}.$
Theorem 5.3
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}.$ Then $\|\mathcal{A}\|_{F}=\frac{1}{\sqrt{n_{3}}}\|\textbf{diag}(\hat{\mathcal{A}})\|_{F}$.
Proof.  By applying identities (3), (12) and $F_{n_{3}}\mathbf{j}=\mathbf{j}F_{n_{3}}$ as well as the invariant property of Frobenius norm under a unitary transformation for complex matrices, we have
$\|\textbf{diag}(\hat{\mathcal{A}})\|_{F}^{2}$ $=\|\textbf{Vec}(\hat{\mathcal{A}})\|_{F}^{2}$
$=\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})+\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\mathbf{j}\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}})\|_{F}^{2}$
$=\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})\|_{F}^{2}+\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}})\|_{F}^{2}$
$=n_{3}\|\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})\|_{F}^{2}+n_{3}\|\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}})\|_{F}^{2}$
$=n_{3}\|\textbf{Vec}(\mathcal{A})\|_{F}^{2}$
$=n_{3}\|\mathcal{A}\|_{F}^{2}.$
Therefore, $\|\mathcal{A}\|_{F}=\frac{1}{\sqrt{n_{3}}}\|\textbf{diag}(\hat{\mathcal{A}})\|_{F}$. $\blacksquare$
Theorem 5.4
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}.$ Then the least-squares solutions of tensor equation (9) can be represented as $\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{I}-\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A})\ast_{Ht}\mathcal{W}$, where $\mathcal{W}$ is an arbitrary compatible RB tensor, and the minimum-norm least-squares solution of tensor equation (9) is $\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}$.
Proof.  According to Theorem 2.3 and Theorem 5.3, we have
$\|\mathcal{A}\ast_{Ht}\mathcal{X}-\mathcal{B}\|_{F}=\frac{1}{\sqrt{n_{3}}}\|\textbf{diag}(\hat{\mathcal{A}})\textbf{diag}(\hat{\mathcal{X}})-\textbf{diag}(\hat{\mathcal{B}})\|_{F}.$
Thus, they arrive the minima at the same time. We know that the the least-squares solutions of RB matrix equation $\textbf{diag}(\hat{\mathcal{A}})\textbf{diag}(\hat{\mathcal{X}})=\textbf{diag}(\hat{\mathcal{B}})$ are
$\textbf{diag}(\hat{\mathcal{X}})=\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{B}})+[\textbf{diag}(\hat{\mathcal{I}})-\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{A}})]\textbf{diag}(\hat{\mathcal{W}})$
with $\textbf{diag}(\hat{\mathcal{W}})$ be arbitrary. By using the fact that $\textbf{diag}(\hat{\mathcal{A}})^{\dagger}=\textbf{diag}(\hat{\mathcal{A}^{\dagger}})$ and Theorem 2.3 we can derive
$\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{I}-\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A})\ast_{Ht}\mathcal{W}$
is the least-squares solution of $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$. For any least-squares solution $\mathcal{X}_{0}$, we have
$\|\mathcal{X}_{0}\|^{2}_{F}=$ $\|\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{I}-\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A})\ast_{Ht}\mathcal{W}\|^{2}_{F}$
$=$ $\frac{1}{n_{3}}\|\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{B}})+(\textbf{diag}(\hat{\mathcal{I}})-\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{A}}))\textbf{diag}(\hat{\mathcal{W}})\|^{2}_{F}$
$=$ $\frac{1}{n_{3}}[\|\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{B}})\|^{2}_{F}+\|\textbf{diag}(\hat{\mathcal{I}})-\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{A}}))\textbf{diag}(\hat{\mathcal{W}})\|^{2}_{F}$
$+2tr(\Re((\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{B}}))^{\ast}(\textbf{diag}(\hat{\mathcal{I}})-\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{A}}))\textbf{diag}(\hat{\mathcal{W}})))].$
Since
$(\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{B}}))^{\ast}(\textbf{diag}(\hat{\mathcal{I}})-\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{A}}))\textbf{diag}(\hat{\mathcal{W}})$
$=$ $\textbf{diag}(\hat{\mathcal{B}}))^{\ast}(\textbf{diag}(\hat{\mathcal{A}})^{\dagger})^{\ast}(\textbf{diag}(\hat{\mathcal{I}})-(\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{A}}))^{\ast})\textbf{diag}(\hat{\mathcal{W}})$
$=$ $\textbf{diag}(\hat{\mathcal{B}}))^{\ast}((\textbf{diag}(\hat{\mathcal{A}})^{\dagger})^{\ast}-(\textbf{diag}(\hat{\mathcal{A}})^{\dagger}\textbf{diag}(\hat{\mathcal{A}})\textbf{diag}(\hat{\mathcal{A}})^{\dagger})^{\ast})\textbf{diag}(\hat{\mathcal{W}})$
$=$ $O,$
we can conclude that
$\|\mathcal{X}_{0}\|^{2}_{F}=\|\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}\|^{2}_{F}+\|(\mathcal{I}-\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A})\ast_{Ht}\mathcal{W}\|^{2}_{F}\geq\|\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}\|^{2}_{F}.$
Consequently, the least-squares solution of tensor equation (9) with minimum norm is $\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}$. $\blacksquare$
In an analogous way, we have the following result. For simplicity, we omit the proof here.
Theorem 5.5
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{4}\times n_{2}\times n_{3}}.$ Then the least-squares solutions of RB tensor equation $\mathcal{X}\ast_{Ht}\mathcal{A}=\mathcal{B}$ can be represented as $\mathcal{X}=\mathcal{B}\ast_{Ht}\mathcal{A}^{\dagger}+\mathcal{W}\ast_{Ht}(\mathcal{I}-\mathcal{A}\ast_{Ht}\mathcal{A}^{\dagger})$, where $\mathcal{W}$ is an arbitrary compatible RB tensor, and the the minimum-norm least-squares solutions of RB tensor equation $\mathcal{X}\ast_{Ht}\mathcal{A}=\mathcal{B}$ is $\mathcal{X}=\mathcal{B}\ast_{Ht}\mathcal{A}^{\dagger}$.
6 Applications in color video processing
In the section, we propose two algorithms for the color video compression and debluring. Both experiments of color video processing show the superiority of our methods over the compared methods.
Based on the Ht-SVD, we define Ht-tubal rank for the tensor $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and the rank-$k$ approximation $\mathcal{A}_{k}$ of $\mathcal{A}$ as follows.
Definition 6.1
For a tensor $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ with Ht-SVD, i.e, $\mathcal{A}=\mathcal{U}*_{Ht}\mathcal{S}*_{Ht}\mathcal{V}^{*}$. The Ht-tubal rank $rank_{Ht-tubal}(\mathcal{A})$ of $\mathcal{A}$ is defined to be the number of nonzero elements of ${\mathcal{S}(i,i,:)}^{w}_{i=1}$, where $w\doteq\text{min}(n_{1},n_{2})$, i.e.,
$rank_{Ht-tubal}(\mathcal{A})=\#\{i\mid\|\mathcal{S}(i,i,:)\|_{F}>0\}.$
When we compute the SVD of frontal slices $\hat{\mathcal{A}}^{(1)},\cdots,\hat{\mathcal{A}}^{(n_{3})}$ with singular values in decreasing order, $rank_{Ht-tubal}(\mathcal{A})=\mathop{\max}\limits_{i}\{rank(\hat{\mathcal{A}}^{(1)}),\cdots,rank(\hat{\mathcal{A}}^{(n_{3})})\}$. If we assume $k\leq rank_{Ht-tubal}(\mathcal{A})$, then
$\mathcal{A}_{k}=\sum\limits_{i=1}^{k}\mathcal{U}(:,i,:)\ast_{Ht}\mathcal{S}(i,i,:)\ast_{Ht}\mathcal{V}(:,i,:)^{\ast}$ (13)
has the Ht-tubal rank of $k.$ Next, we will apply the approximation $\mathcal{A}_{k}$ of $\mathcal{A}$ to compress the color video.
Example 6.2: (Color video compression) We use a reduced biquaternion tensor $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ to express a color video, i.e.,
$\mathcal{A}_{C}(m,n,t)=\mathcal{A}_{R}(m,n,t)\mathbf{i}+\mathcal{A}_{G}(m,n,t)\mathbf{j}+\mathcal{A}_{B}(m,n,t)\mathbf{k},$
where the red $\mathcal{A}_{R}(m,n,t)$, green $\mathcal{A}_{G}(m,n,t)$ and blue $\mathcal{A}_{B}(m,n,t)$ pixel values are sets of the $t$-th color image of a color video.
Algorithm 2 (RkA-RBT) Rank-$k$ Approximation $\mathcal{A}_{k}$ of Reduced Biquaternion Tensor $\mathcal{A}$
1:  Input: $\mathcal{A}_{(c),1},\mathcal{A}_{(c),2}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ from given $\mathcal{A}=\mathcal{A}_{(c),1}e_{1}+\mathcal{A}_{(c),2}e_{2}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$, given $k$
2:  Output: $\mathcal{A}_{k}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\textbf{}\times n_{3}}$ such that
$\mathcal{A}_{k}=\mathcal{U}_{k}\ast_{Ht}\mathcal{S}_{k}\ast_{Ht}\mathcal{V}_{k}^{\ast}$
3:  $\hat{\mathcal{A}}_{(c),1}$=fft$(\mathcal{A}_{(c),1},[],3)$;$\hat{\mathcal{A}}_{(c),2}$=fft$(\mathcal{A}_{(c),2},[],3)$
4:  Calculate each frontal slice of $\hat{\mathcal{U}},\hat{\mathcal{S}}$ and $\hat{\mathcal{V}}$ from $\hat{\mathcal{A}}$:for $i=1,...n_{3},$ do$[\hat{U}_{1}^{(i)},\hat{S}_{1}^{(i)},\hat{V}_{1}^{(i)}]=svd(\hat{\mathcal{A}}_{(c),1}^{(i)})$, $[\hat{U}_{2}^{(i)},\hat{S}_{2}^{(i)},\hat{V}_{2}^{(i)}]=svd(\hat{\mathcal{A}}_{(c),2}^{(i)})$end for
[…]
7:  $\mathcal{A}_{k}=\mathcal{A}_{(c),1}(k)e_{1}+\mathcal{A}_{(c),2}(k)e_{2}$
8:  return $\mathcal{A}_{k}$.
For comparison, we apply SVD-RBT to compress the same color video “DO01_013” as the one in [13], while in [13], it was compressed by Qt-SVD. Fig. 1 shows the PSNRs of the compressed video when taking the rank-$k$ $(k=10,20,50)$ approximations $\mathcal{A}_{k}$ in (13) to “DO01_013” for the 1st, the 20th and the 50th frames, respectively.
To compute the PSNR of each frame, we can treat each frame as a real tensor $\mathcal{C}\in\mathbb{R}^{n_{1}\times n_{2}\times 3}$, and the approximate frame as $\mathcal{C}_{k}\in\mathbb{R}^{n_{1}\times n_{2}\times 3}$, thus its PSNR is defined as
$PSNR\doteq 10log_{10}\left(\frac{3n_{1}n_{2}\|\mathcal{C}\|_{\infty}^{2}}{\|\mathcal{C}-\mathcal{C}_{k}\|_{F}^{2}}\right),$ (14)
with $\mathcal{C}=(C_{mnt})_{n_{1}\times n_{2}\times 3},\mathcal{C}_{k}\in\mathbb{R}^{n_{1}\times n_{2}\times 3}$, $\|\mathcal{C}\|_{\infty}=\mathop{\max}\limits_{m,n,t}|C_{mnt}|$.
From Fig. 1, we see that the performance of RkA-RBT is comparable to Fig. 2 in [13]. Moreover, Table 1 shows that the using of the Ht-SVD is much faster than Qt-SVD in [13], which used the algorithm provided by Jia, Ng and Song [18] in MATLAB.
Figure 1: PSNRs of the rank-k (k=10, 20, 50) approximations to DO01_013 by Ht-SVD.
frame 10 20 30 40
Qt-SVD 20.861257s 46.294602s 89.366687s 116.283031s
Ht-SVD 2.715995s 6.818766s 9.597188s 13.1181154s
Table 1: The running time of Ht-SVD and Qt-SVD for the different number of frames.
Example 6.3: (Color video deblurring) Video deblurring is a crucial task in video processing with applications ranging from video enhancement to computer vision tasks, like object recognition and tracking. Given an original video $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and its blurred video $\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$, the goal is to find the deblurring filter $\mathcal{F}$ that minimizes the reconstruction error between the filtered video $\mathcal{F}\ast_{Ht}\mathcal{B}$ and the original video $\mathcal{A}$.
This problem can be formulated as the following least-squares problem:
$\min_{\mathcal{F}}\|\mathcal{A}-\mathcal{F}\ast_{Ht}\mathcal{B}\|_{F}^{2}.$
Once the deblurring filter $\mathcal{F}$ is obtained, it can be applied to a blurred video to achieve deblurring. Specifically, for a new blurred video $\mathcal{B}_{\text{new}}$, the deblurred video $\mathcal{A}_{\text{deblurred}}$ can be obtained by:
$\mathcal{A}_{\text{deblurred}}=\mathcal{F}\ast_{Ht}\mathcal{B}_{\text{new}}.$
In this problem, the filter $\mathcal{F}$ is determined via a least-squares method. According to Theorem 5.5, the optimal solution for $\mathcal{F}$ is given by $\mathcal{F}=\mathcal{A}\ast_{Ht}\mathcal{B}^{\dagger}$. The detailed solution process is summarized in Algorithm 3.
Algorithm 3 (LSD-RBT) Least-Squares Deblurring via Reduced Biquaternion Tensors
1:  Input: $\mathcal{A}_{(c),1},\mathcal{A}_{(c),2}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ from given $\mathcal{A}=\mathcal{A}_{(c),1}e_{1}+\mathcal{A}_{(c),2}e_{2}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ ; $\mathcal{B}_{(c),1},\mathcal{B}_{(c),2}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ from given $\mathcal{B}=\mathcal{B}_{(c),1}e_{1}+\mathcal{B}_{(c),2}e_{2}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$
2:  Output: $\mathcal{F}\in\mathbb{H}_{c}^{n_{1}\times n_{1}\times n_{3}}$ such that
$\mathcal{F}=\min_{\mathcal{F}}\|\mathcal{A}-\mathcal{F}\ast_{Ht}\mathcal{B}\|_{F}^{2}$
3:  $\hat{\mathcal{A}}_{(c),1}$=fft$(\mathcal{A}_{(c),1},[],3)$;$\hat{\mathcal{A}}_{(c),2}$=fft$(\mathcal{A}_{(c),2},[],3);\hat{\mathcal{B}}_{(c),1}$=fft$(\mathcal{B}_{(c),1},[],3)$;^$\hat{\mathcal{B}}_{(c),2}$=fft$(\mathcal{B}_{(c),2},[],3)$
4:  for $i=1,...n_{3},$ do$[\hat{U}_{1}^{(i)},\hat{S}_{1}^{(i)},\hat{V}_{1}^{(i)}]=svd(\hat{\mathcal{B}}_{(c),1}^{(i)})$, $[\hat{U}_{2}^{(i)},\hat{S}_{2}^{(i)},\hat{V}_{2}^{(i)}]=svd(\hat{\mathcal{B}}_{(c),2}^{(i)})$^$\hat{\mathcal{B}}_{(c),1}^{\dagger(i)}=\hat{V}_{1}^{(i)}\hat{S}_{1}^{\dagger(i)}\hat{U}_{1}^{(i)\ast}$, ^$\hat{\mathcal{B}}_{(c),2}^{\dagger(i)}=\hat{V}_{2}^{(i)}\hat{S}_{2}^{\dagger(i)}\hat{U}_{2}^{(i)\ast}$$\hat{\mathcal{F}}_{(c),1}=\hat{\mathcal{A}}_{(c),1}^{(i)}\hat{\mathcal{B}}_{(c),1}^{\dagger(i)}$, ^$\hat{\mathcal{F}}_{(c),2}=\hat{\mathcal{A}}_{(c),2}^{(i)}\hat{\mathcal{B}}_{(c),2}^{\dagger(i)}$ end for
[…]
6:  return $\mathcal{F}$
For video deblurring experiments, we select color videos from the DAVIS 2017 dataset [17]. The experimental comparison is conducted between traditional quaternion-based method [19] and the newly proposed reduced biquaternion method. Table 2 illustrates the comparative results in terms of PSNR, relative error, and CPU time per frame. The relative error between the deblurred video $\tilde{\mathcal{A}}$ and the original video $\mathcal{A}$ is computed as follows:
Relative $\text{Relative Error}=\frac{\|\mathcal{A}-\tilde{\mathcal{A}}\|_{F}}{\|\mathcal{A}\|_{F}},$
and PSNR is defined by (14). As shown in Table 2, our reduced biquaternion method LSD-RDT demonstrates superior performance compared to traditional quaternion-based method. Fig. 2 and Fig. 3 showcase the enhanced deblurring effects on the ’Tram’ and ’Butterfly’ videos, respectively, illustrating the effectiveness of the RB method in restoring video clarity.
Video Method PSNR Relative error CPU time (second)
Tram Quaternion 30.4592 0.058439 1.867
Reduced biquaternion 33.8519 0.039519 0.152
Butterfly Quaternion 32.9166 0.053911 1.947
Reduced biquaternion 38.8111 0.027293 0.147
Table 2: Performance Comparison of Quaternion and Reduced Biquaternion Methods for Video Deblurring.
Figure 2: Example of Deblurring Estimation for Three Blurred Frames in Video ’Tram’.
Figure 3: Example of Deblurring Estimation for Three Blurred Frames in Video ’Butterfly’.
Conclusion: In this paper, we first define Ht-product for the third-order RB tensors. Then under this product, we define the Ht-SVD for a third-order RB tensor and develop a computing algorithm. Moreover, we extend the Moore-Penrose inverse to the third-order RB tensors. Our Ht-SVD allows us to prove many properties of this MP inverse efficiently. Furthermore, we discuss the general solution/the least-squares solutions/minimal norm least-squares solution to the classic RB tensor equations via our MP inverses. As the applications of our results, we develop two algorithms and apply them to color video compression and deblurring. Both of two experiments illustrate the superiority of our methods by comparing with other methods.
7 Declarations
Conflict of Interest: The authors have not disclosed any competing interests.
Data Availability Statement: The data supporting the findings of this study are not publicly available due to privacy or ethical restrictions. However, interested researchers may request access to the data by contacting the corresponding author and completing any necessary data sharing agreements.